### v1

# IJCAI–25 Formatting Instructions

Abstract
Conventional retrieval-augmented neural machine translation (RANMT) systems leverage bilingual corpora, e.g., translation memories (TMs). Yet, in many settings, in-domain monolingual target-side corpora are often available. This work explores ways to take advantage of such resources by retrieving relevant segments directly in the target language, based on a source-side query. For this, we design improved cross-lingual retrieval systems, trained with both sentence level and word-level matching objectives. In our experiments with two RANMT architectures, we first demonstrate the benefits of such cross-lingual objectives in a controlled setting, obtaining translation performances that surpass standard TM-based models. We then showcase our method on a real-world set-up, where the target monolingual resources far exceed the amount of parallel data and observe large improvements of our new techniques, which outperform both the baseline setting, and general-purpose cross-lingual retrievers.
1 Introduction
The use of Retrieval-Augmented Generative Models (Li et al., 2022) is rapidly expanding, owing to their built-in ability to condition generations with illustrative segments retrieved from memory. The combination of (machine) retrieval and (human) regeneration has long been a key principle of Computer Aided Translation (CAT) systems (Arthern, 1978; Kay, 1997; Bowker, 2002): segments resembling the source sentence are first retrieved from Translation Memories (TM), providing translators valuable suggestions that can then be edited into a new translation. Such techniques have been transposed in Statistical MT (Koehn and Senellart, 2010), more recently in Retrieval-Augmented Neural Machine Translation (RANMT) e.g., (Gu et al., 2018), making the entire process fully automatic. Most subsequent work has continued to leverage parallel data in the form of TMs, performing retrieval in the source language, then using the retrieved target side text in generation. Similar principles underlie few-shot MT with large language models, where both the source and target sides of examples are inserted into the prompt Brown et al. (2020).
TM-based approaches typically retrieve examples with lexical fuzzy matchers, using BM25 and/or the Levenshtein distance (LED) as filters or rankers (Bouthors et al., 2024). These are known to be computationally efficient, and to often surpass continuous-space retrievers (Xu et al., 2020): this is because lexical (source) matches often come with appropriate translation of terms. However, relying on source-side lexical similarity raises two issues that can cause irrelevant segments to be retrieved: (a) noisy alignments between source and target side of a parallel instance; (b) intrinsic morphosyntactic divergences between the two languages (Dorr, 1994).
These problems can be sidestep by performing retrieval directly on the target side, using Cross-Lingual Information Retrieval (CLIR) techniques. An important additional benefit would be to dispense with the use of parallel data and to potentially inform the generation process with relevant monolingual resources. As demonstrated by the recurrent use of target side monolingual data thanks to back-translation Sennrich et al. (2016), such resources are much easier to find than parallel ones in many domains, and also devoid of so-called translationese phenomena Bogoychev and Sennrich (2020). As such, monolingual corpus constitute a very useful, yet under-studied, source of high-quality segments for RANMT. To search such resources, cross-lingual retrievers are readily available thanks to multilingual sentence encoders such as LASER (Artetxe and Schwenk, 2019b) or LaBSE (Feng et al., 2022). They however lack the critical ability to ensure that the retrieved examples will have a large lexical overlap with the source text.
In this work, we propose novel methods to boost the lexical matching abilities of language-agnostic sentence encoders, thereby improving their effectiveness for RANMT. For this, we consider various ways to fine-tune a pre-trained encoder into reproducing the behaviour of lexical fuzzy matchers. We experiment with two language pairs (English-French and German-English) and multiple domains of various sizes and diversities. We consider two types of RANMT systems: an autoregressive source-augmented transformer (Bulte and Tezcan, 2019) and a non-autoregressive edit-based model (Bouthors et al., 2023). We first study the performance of CLIR techniques in a controlled setting, showing that they match standard TM-based models with source-side retrieval. We then consider a much larger set-up, where the target monolingual resources far exceed the amount of parallel data and observe large improvements using our new techniques, which outperform both the baseline setting, as well as generic cross-lingual retrievers.
2 Related Work
Our work falls under the scope of Retrieval-Augmented Generation, leveraging both lexical (BM25 (Robertson and Walker, 1994), Levenshtein distance (Levenshtein, 1965)) and semantic methods for the Information Retrieval (IR) component. Semantic methods usually rely on dual encoders trained on parallel data (Gillick et al., 2018) with a contrastive loss (Sohn, 2016). Our focus is on massively multilingual sentence encoders that can encode mutual translations as close neighbors, such as e.g, LASER (Artetxe and Schwenk, 2019b) and LaBSE (Feng et al., 2022).
In recent years, there has been a growing focus on integrating retrieval components into NMT systems. One key motivation is to enhance model transparency by providing users with the retrieved examples that inform the generated output (Rudin, 2019). A simple implementation of this idea encodes the target side of the example(s) along with the source sentence, providing an additional translation context (Bulte and Tezcan, 2019; Gu et al., 2018; Xia et al., 2019; He et al., 2021; Cheng et al., 2022; Agrawal et al., 2023). It is also possible to leverage both the source and target sides of the example(s) (Pham et al., 2020; Reheman et al., 2023).
Instead of regenerating a complete text with an extended context, edit-based models aim to perform minimal changes to the retrived examples and generate their output in a non-autoregressive (NAR) fashion. Such techniques are studied in (Niwa et al., 2022; Xu et al., 2023; Zheng et al., 2023), adapting the Levenshtein transformer (Gu et al., 2019). This approach is extanded to handle multiple examples that are simultaneously edited in (Bouthors et al., 2023).
MT systems based on Large Language Models (LLMs) can also seamlessly accomodate examples though In-Context Learning (ICL), where the LLM prompt is enriched with multiple demonstrations of the translation task, comprising both the source and target sides of parallel sentences (Radford et al., 2019). Several studies have focused on optimizing the performance of such architectures for NMT, examining the impact of prompt modifications, the number of demonstrations and the retrieval procedure (Moslem et al., 2023; Vilar et al., 2023; Zhang et al., 2023; Hendy et al., 2023; Bawden and Yvon, 2023; Zebaze et al., 2024).
Our lexical alignment objective corresponds to a form of neural metric learning, where a high dimensional distance space is reduced to a smaller one (Suárez et al., 2021). This field of research aims to learn distances from the data (Kulis, 2013; Bellet et al., 2015), which is particulary useful for similarity-based algorithms used in IR (Cakir et al., 2019).
3 Method
3.1 Problem Formulation
Given a source sentence $\mathbf{x}$ and a corpus of segments $\mathcal{D}=(\mathbf{y}_{1},\dots,\mathbf{y}_{N})$ in the target language, the goal of CLIR in NMT is to retrieve $k$ relevant examples: (~$(\tilde{}\mathbf{y}_{1},\dots,\tilde{}\mathbf{y}_{k})$.
Our approach closely follows neural-based cross-lingual retrievers, and comprises two main steps. First, $\mathcal{D}$ is encoded with a multilingual sentence encoder $E_{\theta}$. Then, for any query source sentence $\mathbf{x}$, retrieval is performed as a $k$NN similarity search in the embedding space, using the cosine similarity:
$\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y})=\cos(E_{\theta}(\mathbf{x}),E_{\theta}(\tilde{}\mathbf{y})).$ (1)
The quality of the retrieved sentences only depends on the choice of $E_{\theta}$. Usually, $E_{\theta}$ is trained as a Siamese encoder with a contrastive loss on parallel data, a usual IR objective (Sohn, 2016) that primarily captures semantic similarity. However, it is desirable for the retrieved $\tilde{\mathbf{y}}_{i}$ instances to exhibit surface-level similarities to what could potentially be a translation $\mathbf{y}$ of the source $\mathbf{x}$. In specialized domains (like law or medicine), adherence to domain-specific terminology takes precedence over semantic similarity, thereby motivating the design of our proposed lexical-based alignment objective. To address this need, we study various fine-tuning objectives based on the Levenshtein similarity. These objectives are specifically designed to boost the retrieval of segments with a large numbers of lexical matches in the source query, by learning a new embedding space, achieved through the optimization of parameters $\theta$, such that:
$f(\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}))\approx\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}),$ (2)
where $f$ is a (potentially learnt) mapping from the codomain of the cosine function into the codomain of Levenshtein similarity function $\operatorname{Lev}$, defined as:
$\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})=1-\frac{\Delta(\mathbf{y},\tilde{}\mathbf{y})}{\max(|\mathbf{y}|,|\tilde{}\mathbf{y}|)},$ (3)
where $\Delta$ is the Levenshtein distance (Levenshtein, 1965).
3.2 Training
We consider three objective functions. The first two learn to predict the Levenshtein similarity. We choose $f$ to be a learnable fonction that maps $[-1,1]$ to $[0,1]$:
$f(t)=\sigma(a\times\operatorname{arctanh}(t)+b),$ (4)
with $a$ (slope) and $b$ (position) parameters and $\sigma$ the sigmoid function. The loss is defined as:
$\mathcal{L}(\mathbf{x},\mathbf{y},\tilde{}\mathbf{y})=\operatorname{Err}(f(\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y})),\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})),$ (5)
where the error function $\operatorname{Err}$ is either the mean square error (MSE) or the mean absolute error (MAE) defined as:
$\operatorname{MSE}(x,y)=(x-y)^{2}$ (6)
$\operatorname{MAE}(x,y)=|x-y|.$ (7)
The third objective is a “learning to rank” objective (Cao et al., 2007). Given a set of $k$ retrieved segments ~$\tilde{}\mathbf{y}_{[1:k]}$ sorted in decreasing order w.r.t. $\operatorname{Lev}(\mathbf{y},\cdot)$, it computes:
$\mathcal{L}(\mathbf{x},\mathbf{y},\tilde{}\mathbf{y}_{[1:k]})=$ $\sum_{i>j}\max(0,\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{i})-\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{j})+$
$m\times|\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})-\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})|)$ (8)
with $m$ a scaling factor. Equation (8) defines a pairwise margin loss, [^1] Derived from the triplet loss (Schultz and Joachims, 2003). with an adaptive margin proportional to the absolute difference between the two individual $\operatorname{Lev}$ similarities.
4 Data and Metrics
4.1 Data
We consider two translation tasks: from English to French (en-fr) and from German to English (de-en). Each test corpus contains a wide variety of textual domains (16), all from the Opus corpus (Tiedemann, 2012). The details are shown in appendix A.
We also prepare a large corpus of Wikipedia pages in French. [^2] https://huggingface.co/datasets/Plim/fr_wikipedia_processed. We split the text in pseudo-sentences via regular expression matching and remove duplicates as well as fuzzy matches from the valid/test sets [^3] Each best match with $\operatorname{Lev}>0.9$ is simply removed. of the parallel Wikipedia corpus split, ending up with about $45$M French segments.
4.2 Metrics
The success of cross-lingual retrieval is assessed with the average target side Levenshtein similarity (eq. (3)) between the 1-best match and the reference, for sentences in the validation set, with retrieval performed from the entire train set. We additionally report xsim error rates, which measure the abililty to identify matched sentences in a bilingual parallel corpus (Artetxe and Schwenk, 2019a).
We assess translation quality with BLEU scores (Papineni et al., 2002) computed with SacreBLEU (Post, 2018),444signature: nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.1.0; and COMET-20555Unbabel/wmt22-comet-da, using the official implementation. scores (Rei et al., 2020).
5 Experimental Settings
5.1 Retrieval Settings
Our controlled experiments use parallel corpora in two language pairs (en-fr and de-en). This allows us to perform a head-to-head comparison between standard source-side fuzzy matching and our novel cross-lingual retrieval methods. In each configuration, we retrieve up to $k=3$ closest examples. We compare the following retrieval methods:
  • fuzzy-src: source-to-source fuzzy matching ranked by their $\operatorname{Lev}$ similarity, as in standard RANMT settings;
  • fuzzy-gold: target-to-target fuzzy matching, using the same procedure as fuzzy-src. This setup requires access to reference target translations and corresponds to an oracle score, giving an empirical upper-bound of what an ideal retriever could achieve;
  • fuzzy-bt: the target side of training segments are back-translated into the source language Sennrich et al. (2016), then used for retrieval as in fuzzy-src. This simulates an alternative to cross-lingual retrieval in situations where only target side examples are available;
  • dense: a BERT base Devlin et al. (2019) model trained from scratch with a contrastive loss (Sohn, 2016), usable for source-to-target retrieval. Given a batch $\mathcal{B}=((\mathbf{x}_{1},\mathbf{y}_{1}),\dots,(\mathbf{x}_{n},\mathbf{y}_{n}))$, the loss is computed as:
    $\mathcal{L}(\mathcal{B})=-\sum_{i}\log\frac{\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{i}))}{\sum_{j}\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{j}))};$ (9)
  • dense+bow: we add a bag-of-word loss to dense as in (Cai et al., 2021) to enforce lexical information in cross-lingual encodings. Given a pair of parallel sentences $(\mathbf{x},\mathbf{y})$, the loss is computed as:
    $\mathcal{L}(\mathbf{x},\mathbf{y})=-\sum_{w\in\mathcal{Y}}\log p_{\theta_{1}}(w|\mathbf{x})-\sum_{w\in\mathcal{X}}\log p_{\theta_{2}}(w|\mathbf{y}),$ (10)
    where $\mathcal{X}$, $\mathcal{Y}$ are the bag-of-words of $\mathbf{x}$ et $\mathbf{y}$; $p_{\theta_{1}}$ and $p_{\theta_{2}}$ are distinct linear projections of $E_{\theta}(\mathbf{x})$ and $E_{\theta}(\mathbf{y})$;
  • LASER: we use LASER (Artetxe and Schwenk, 2019b) as a first pre-trained source-to-target retriever. LASER derives from a multilingual RNN-based translation model, repurposing the source-side embeddings to define a similarity measure between multilingual sentence representations;
  • LaBSE: Similar to dense, LaBSE (Feng et al., 2022) uses a dual encoder model based on the BERT base architecture and is trained with a contrastive loss; it additionnally benefits from pre-training on large sets of monolingual data, and covers more than 110 languages;
[…]
In our setting, fine-tuning updates all the model parameters. The global learning rate is 1e-4, as for $a$ and $b$ (eq. 4) it is set to 1e-2. Each source $\mathbf{x}$ is presented with three ~$\tilde{}\mathbf{y}$ (eq. 5) determined by the top 3 scoring matches obtained with dense+bow. The validation score is an in-batch normalized discounted cumulative gain (NDCG) score (Wang et al., 2013).
The retrieval is always performed in an in-domain setting. For all methods, a threshold is further applied to filter out low similarity segments. For each retrieval method, we adjust this threshold on the validation set to achieve an averaged retrieval rate [^6] Proportion of sentences with at least one retrieved example. of $50$%. Using a constant retrieval rate for all retrievers ensures that no system is unduly rewarded for retrieving a greater number of similar examples, and that our comparisons of retrievers only reflects differences in the quality of the TM matches. Please, refer to appendix B for the complete search configurations.
5.2 NMT Architectures
We consider two RANMT architectures. The first is an in-house implementation of NFA (Bulte and Tezcan, 2019), which extends a basic encoder-decoder MT architecture by prepending the target side of TM examples to the source text on the encoder side. As is standard, the retrieved examples used in training are selected based on their Levenshtein distance to the source text; at most one similar example is used. This model thus also has the ability to perform inference from scratch, without using any retrieved example.
The second architecture implements an edit-based, non-autoregressive encoder-decoder model that has the ability to jointly edit up to $k$ examples to generate a translation, reproducing the multi-Levenshtein transformer (TMk-LevT) (Bouthors et al., 2023). Retrieved examples used in training are also selected based on source-side similarity measures; we use up to $k=3$ examples.
6 Results
We first report the results of controlled experiments, in order to contrast mono-lingual vs. cross-lingual retrieval strategies. We then analyze the results of a larger-scale setting, where we can only exploit monolingual data.
test en-fr test de-en
$\operatorname{Lev}$uparrow xsim $\downarrow$ $\operatorname{Lev}$uparrow xsim $\downarrow$
fuzzy-gold (oracle) 50.0 - 53.4 -
fuzzy-src 36.0 - 21.9 -
fuzzy-bt 31.1 - 21.8 -
dense 30.4 10.8 - -
dense+bow 32.7 8.0 30.9 25.2
ft-dense+bow-MSE 34.9 7.6 33.6 25.2
ft-dense+bow-MAE 35.2 8.0 32.9 25.3
ft-dense+bow-Rank 29.9 11.9 25.9 41.5
LASER 32.8 10.0 33.2 26.2
LaBSE 33.8 7.4 34.4 21.0
ft-LaBSE-MSE 36.7 7.6 34.0 24.6
ft-LaBSE-MAE 37.0 6.4 34.5 23.5
ft-LaBSE-Rank 36.0 7.3 34.2 23.3
Table 1: Average retrieval scores (xsim error rate and average Levenshtein similarity), averaged for all domains.
6.1 Retrieval Scores
Table 1 stores the retrieval scores (§4.2) for the small-scale experiments. For en-fr, source-side fuzzy matching yields more similar segments than CLIR techniques, a gap that vanishes after fine-tuning with MAE for dense+bow and even more so for LaBSE. For de-en, the source-side fuzzy matching baseline is much worse, revealing a large gap with the oracle condition, which might be due to residual misalignments or to morphological differences between German and English. [^7] Inflection phenomena in German create multiple variants for each lexeme, making exact matches at the word level much less likely than in English. All CLIR techniques widely outperform the monolingual settings (fuzzy-src and fuzzy-bt). Here, however we do not see much gain from fine-tuning, perhaps owing to a smaller fune-tuning dataset. As expected, fine-tuning with a lexical loss is slightly detrimental to the sentence-level retrieval scores (xsim error) for both dense+bow and LaBSE. Training dense+bow along with a bag-of-word loss (eq. (10)) effectively improves the retrieval scores.
6.2 Translation Scores
Retrieval-free Baselines
We use both NLLB and NFA without any retrieved examples as our machine translation baselines. Their corresponding BLEU scores, averaged across domains, are respectively 39.0 (en-fr) and 33.8 (de-en) for NLLB and 48.0 (en-fr) and 38.2 (de-en) for NFA. Details per domain are in Table 4.
TM3-LevT NFA
test en-fr test de-en test en-fr test de-en
BLEU COMET BLEU COMET BLEU COMET BLEU COMET
sans exemple - - - - 48.0 87.1 38.2 79.8
fuzzy-gold (oracle) 45,3 51,7 31,3 -8,6 52,2 87,8 44,7 82,2
fuzzy-src 43,8 48,4 28,8 -10,3 51,3 87,5 41,9 81,7
fuzzy-bt 40,5 43,8 22,0 -22,9 49,9 87,2 38,9 81,3
dense 41,5 42,2 - - 49,9 87,1 - -
dense+bow 42,4 44,1 28,3 -11,6 50,5 87,3 40,0 80,9
ft-dense+bow-MSE 43,1 46,0 29,0 -12,7 51,0 87,5 41,6 81,5
ft-dense+bow-MAE 43,2 46,3 29,1 -12,6 51,0 87,4 41,8 81,4
ft-dense+bow-Rank 42,1 44,3 26,0 -17,9 50,7 87,3 41,8 80,9
LASER 42,5 44,3 30,4 -7,7 49,9 87,3 41,1 81,3
LaBSE 43,1 45,5 29,8 -9,5 50,5 87,3 41,5 81,4
ft-LaBSE-MSE 43,7 47,8 27,7 -12,5 51,2 87,5 41,6 81,6
ft-LaBSE-MAE 43,6 47,7 29,1 -9,2 51,2 87,5 42,0 81,5
ft-LaBSE-Rank 43,3 47,1 28,4 -12,2 51,1 87,5 41,8 81,6
Table 2: Average translation scores: TM3-LevT and NFA, all domains.
English-French German-English
ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
fuzzy-gold 58.5 64.8 32.5 61.3 60.5 44.3 27.0 41.0 31.2 40.5 36.9 32.1 8.2 48.7 47.6 19.8
fuzzy-src 56.7 62.2 31.9 58.9 58.8 41.5 27.2 40.3 30.7 38.9 34.6 29.2 8.7 45.8 43.3 16.9
fuzzy-bt 51.0 54.7 31.9 51.3 55.1 35.2 27.0 38.7 30.7 36.2 33.3 21.3 4.5 27.5 39.2 17.5
LASER 56.5 60.5 31.7 56.0 58.6 36.6 26.9 39.7 30.6 36.4 33.9 29.3 11.3 48.9 46.6 16.2
LaBSE 57.0 59.8 31.3 57.8 59.4 40.8 26.9 39.8 30.5 37.0 33.9 29.0 9.3 49.2 45.4 16.2
ft-LaBSE-MSE 57.0 62.0 31.8 58.6 58.6 42.5 27.0 39.9 30.5 38.1 34.9 28.1 10.9 41.4 41.1 17.0
ft-LaBSE-MAE 56.9 61.4 31.6 58.7 58.7 42.5 27.0 39.9 30.4 37.7 34.5 28.7 11.7 44.4 43.7 17.0
ft-LaBSE-Rank 56.6 61.1 31.7 57.7 58.0 41.7 26.9 40.1 30.6 37.6 34.1 28.7 9.6 44.6 42.5 16.6
Table 3: Per domain BLEU scores for TM3-LevT systems. Contaminated domains for NLLB (used in back-translation) are underlined.
English-French German-English
ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
contam. rate 33.9 58.3 8.9 0 2.1 6.8 100 6.8 29.8 0 0 47.5 0 54.0 17.3 8.9
NLLB 44.5 40.3 35.9 39.6 50.7 34.2 32.1 39.0 41.0 35.9 35.4 29.1 23.5 46.4 38.8 31.4
no example 58.9 55.9 39.9 48.8 62.3 42.8 50.1 45.6 43.9 43.5 36.4 39.5 13.2 55.5 51.6 31.0
fuzzy-gold 65.2 65.1 40.3 59.7 68.1 46.6 50.1 48.1 44.1 47.8 39.0 45.6 21.9 63.9 60.5 31.7
fuzzy-src 64.5 63.0 40.0 58.4 66.6 45.2 50.0 47.3 44.0 46.8 38.3 44.3 14.2 62.3 57.4 31.1
fuzzy-bt 62.4 60.4 39.8 54.7 64.9 44.5 50.1 46.8 43.7 44.7 37.0 40.1 13.2 55.6 54.8 31.1
LASER 63.7 61.6 39.8 56.8 65.4 44.0 50.0 46.8 44.1 46.1 37.1 41.4 16.1 60.7 56.0 31.1
LaBSE 63.8 62.5 39.0 57.7 65.0 45.3 50.0 46.9 44.1 46.0 37.7 43.0 15.6 60.6 56.9 31.1
ft-LaBSE-MSE 64.3 63.0 40.0 58.2 66.6 45.3 50.1 47.5 44.0 46.8 37.9 43.1 16.8 60.3 56.9 31.2
ft-LaBSE-MAE 64.2 63.0 39.9 58.5 66.6 45.5 50.1 47.4 43.9 46.3 37.8 43.3 17.1 61.4 57.5 30.9
ft-LaBSE-Rank 64.1 62.6 39.8 58.1 66.3 45.3 50.0 47.5 44.0 46.0 38.1 43.7 15.9 61.0 57.2 31.2
Table 4: Per domain BLEU scores for NFA systems and NLLB. The contamination rate (top line) for NFA is the percentage of overlap of its training data with the test set. Contaminated domains for NLLB (used in back-translation) are underlined.
TM3-LevT
Translation results achieved with TM3-LevT are in Table 2. The difference between the oracle (fuzzy-gold) and the baseline (fuzzy-src) gives an idea of the quality of the monolingual matches: the quality is pretty high in en-fr, owing to the cleaning procedure used for these data, more noisy in de-en, with a gap of about $2.5$ BLEU points. In comparison, using artificial back-translated source queries performs much worse than using natural texts. This difference is much larger than what was expected from the retrieval scores of Table 1.
For both language pairs, using cross-lingual search enables to obtain results that compare to source-side fuzzy matching. For en-fr, the best scores in our setting use fine-tuned versions of LaBSE, closing the gap in BLEU with fuzzy-src, and delivering a small COMET gain. Similar to the retrieval scores (table 1), we do not observe such benefit of fine-tuning for de-en, where cross-lingual retrieval with LASER yields the best overall results, yielding a $1.6$ BLEU points improvement over the pure source-side matching.
NFA
The corresponding results for NFA are in Table 2. This model is a much stronger baseline than TM3-LevT and the margin of progress (up to fuzzy-gold) is dimmer. Yet, we see that (a) retrieving examples directly on the target side (with dense, LASER, and LaBSE) already improves the example-free baseline by about $3$ BLEU points; (b) an extra step of fine-tuning of cross-lingual retrievers (with MSE and MAE) enables us to fully close the gap with source-side matching for both metrics and language pairs. In comparison, back-translating the monolingual target data also improves our baseline systems, but by a much smaller margin.
Per-domain analysis
Per-domain BLEU scores are in Tables 3 (TM3-LevT) and 4 (NFA). The domains for which we expect CLIR methods to work well need to satisfy two conditions: (a) a large enough set of examples to retrieve from, so that the choice the retriever actually makes a difference; (b) a small diversity of texts in that domain, so that retrieved instances are sufficiently similar to the input. This is well reflected in Table 3 (TM3-LevT) where CLIR techniques (with fine-tuning in en-fr, without in de-en) match the performance of the baseline in all domains, and even outpeform it in specialized domains such as ECB (finance), JRC (law), KDE (information technology) and Wikipedia. We do not see the same gains in Ubuntu (very small TM) or Europarl (large TM, with a diverse set of texts).
For NFA, we also observe that using CLIR techniques yields scores that match the baseline; for both monolingual or cross-lingual retrieval, the improvement margin, as defined by the difference with fuzzy-gold, is small for all domains, except for Koran (de-en), where CLIR techniques once again demonstrate their effectiveness.
Comparison of fine-tuning losses
The standard contrastive loss of equation (9), when applied in MT, primarily seeks to identify parallel sentences within collections of monolingual segments. Yet, RANMT requires examples that are lexically similar to the source sentences. The bag-of-word loss of equation (10) effectively increases the rate of lexically similar sentences, thus enhancing the translation scores. As for our three fine-tuning strategies, MSE and MAE (eq. (5)) yield comparable performances. The per-domain analysis further highlights their consistent behavior. As for Rank (eq. (8)), we observe that the rank-based loss consistently underperforms the other two, with a few exceptions (PHP for NFA; PHP, Wiki and KDE (de-en) for TM3-LevT).
On data contamination
We identified two sources of data contamination. First, NFA training data partly overlaps with some of our test sets, in varying proportion (from 0 to 100) across domains (see Table 4). The intuition is that such contamination should make this model less sensitive to the choice of good examples. In the per-domain results of table 4, this sensitivity is assessed by the gap between example-free and fuzzy-gold scores. As this gap is large for most domains (with the exception of Europarl and News in en-fr, and Sub in de-en), we can still clearly assess the benefits of cross-lingual retrieval for the large majority of domains. Second, the training data of NLLB, used for back-translation, also partly overlaps with our test, which may boost the efficiency of the fuzzy-bt baseline. Even for these domains (e.g., ECB, EMEA, JRC-Acquis), cross-lingual methods still achieve much higher BLEU scores than back-translation.
Figure 1: Evolution of BLEU and COMET scores of TM3-LevT (top) and NFA (bottom) when growing the retrieval pool.
6.3 A Large-scale Experiment
Retrieving in a monolingual corpus
We use the models trained on the en-fr dataset to retrieve segments from the large monolingual Wikipedia dataset (see §4.1), using the Wikipedia test already used in the controled experiments set to assess performance. We use the same filtering threshold as for the controlled experiments to keep the settings comparable and ensure that the only difference is the increased size of the retriveal set. We only consider the variants of LaBSE, which achieve the best results in the small data condition.
Wikipedia is a challenging domain, due to the variety of topics that are addressed in the encyclopedia and the high level of formality of the text. NLLB already represents a strong baseline for that domain. Results are in Table 5. The first observation is that increasing the retrieval set (from 200K segments in the controlled experiments – Tables 3 and 4 – to 45M segments) yields significant performance boosts across the board (e.g, +1.6 BLEU for NFA and fuzzy-bt, +0.4 BLEU points for NFA and LaBSE). With an additional lexical fine-tuning step, the benefits of the cross-lingual approach are again very clear, with large BLEU differences relative to the small data condition:+3.9 for ft-LaBSE-MAE with TM3-LevT, +3.4 when using NFA.
TM3-LevT NFA
BLEU COMET BLEU COMET
no example - - 36.4 84.8
fuzzy-bt 35.6 76.7 38.6 84.8
LaBSE 37.7 77.2 40.1 85.0
ft-LaBSE-MSE 38.3 77.2 40.9 84.8
ft-LaBSE-MAE 38.7 77.6 41.2 84.9
ft-LaBSE-Rank 38.5 77.5 40.8 85.0
Table 5: BLEU and COMET scores with systems TM3-LevT and NFA with the Wikipedia retrieval pool.
Effect of the retrieval pool size
To better analyse the effect of the retrieval pool size, we progressively increase the Wikipedia retrieval set by batches of $500$K sentences from $500$K to $45$M, and compute the corresponding translation scores for each of the retrieval systems. Figure 1 reports the evolution of the translation metrics for LaBSE, its fine-tuned variants and fuzzy-bt. Overall, for all models, the benefit of growing the retrieval pool yields an almost linear BLEU increase. COMET variations are much smaller, especially for NFA where it remains nearly constant for our best setups. For both metrics, ft-LaBSE-MAE outperforms the other setups for all data sizes. fuzzy-bt consistently lags behind CLIR alternatives, especially with TM3-LevT.
We also monitor in table 6 two additional scores: the retrieval rate (percentage of segments for which at least one example is found) and the average $\operatorname{Lev}$ similarity between the reference and its closest retrieved neighbor. For all retrieval methods, we see large increases, meaning a larger number of retrieved examples is retrieved, and that their average quality (i.e their similarity to the reference) also improves, highliting the usefulness of monolingual data for this domain.
TM mono
RR $\uparrow$ $\operatorname{Lev}$uparrow RR $\uparrow$ $\operatorname{Lev}$uparrow
fuzzy-bt 39.9 24.0 60.2 31.0
LaBSE 30.9 25.6 59.6 35.2
ft-LaBSE-MSE 33.8 27.2 77.7 37.7
ft-LaBSE-MAE 31.3 27.5 59.0 38.1
ft-LaBSE-Rank 25.7 26.7 61.2 36.7
Table 6: Comparison of the retrieval rates (RR) and average Levenshtein similarity ($\operatorname{Lev}$) on the Wikipedia test set of the systems retrieving in the TM vs. in the large monolingual corpus (mono).
7 Conclusions and Outlook
In this paper, we have explored various approaches to take advantage of monolingual examples in retrieval-augmented neural machine translation. Our main conclusion, resulting from experiments on 2 language directions and 16 varied textual domains, is that retrieving directly in the target language can be as effective as fuzzy-matching on the source-side, especially with lexical-based fine-tuning objectives. Using these techniques, we were also able to obtain large BLEU increases over strong baselines on the Wikipedia dataset (up to +3.8 BLEU points for our best configuration), showcasing the benefits of RANMT in a large-scale setup. In our implementation, cross-lingual retrieval is no more costly than performing fuzzy matches in the source language.
We identify several ways to improve these results: (a) by fine-tuning the retrieval models on very large multilingual datasets, adopting data settings that are comparable to the pre-training of LASER/LaBSE; (b) using more recent sentence embedders such as SONAR (Duquenne et al., 2023); (c) using CLIR techniques to train the RANMT models, instead of relying on source-side matches as we have done here; (d) revisiting the other aspect of the retrieval pipeline – e.g., relaxing the filtering threshold during example selection. Another continuation of this work is to adopt CLIR for in-context learning with LLMs, which are currently instruction-tuned with parallel examples, based on source-side similarities Moslem et al. (2023). Fine-tuning LLMs with just target side examples retrieved with CLIR is one priority of our future work. Admittedly, we worked with high ressource languages to assess the viability of our method. Considering a low ressource setup, provided large enough monolingual corpora are available, is a natural use case when it is even harder to access to parallel data. Such experiments can further assess the strenght of our method.
Appendix A Parallel Data Specification
English-French German-English
ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
size 149k 272k 1,9M 44k 492k 126k 133k 7k 148k 6k 597k 223k 18k 467k 248k 500k
% 3.8 7.0 49.0 1.1 12.7 3.3 3.4 0.2 3.8 0.2 15.4 15,3 1,2 32,1 17,0 34,3
Table 7: Size of each domain, and its proportion w.r.t. the size of the corpus of the same language pair.
The en-fr data is a clean version [^8] We filter noisy parallel pairs with COMETKiwi Rei et al. [2022] and prepare a new train/valid/test split. of the one used by [Xu et al., 2020]. Validation and test sets each contain $1,000$ segments. The de-en corpus is borrowed unchanged from [Aharoni and Goldberg, 2020] and is also used in [Cai et al., 2021; Agrawal et al., 2023; Bouthors et al., 2024]. The associated validation and test sets contain $2,000$ segments.
Appendix B Search Parameterization
For both methods, we search up to $k$=3 examples.
B.1 Fuzzy Matching
The execution of the fuzzy matching methods (fuzzy-src, fuzzy-bt and fuzzy-gold) is done via an open source library [^9] https://github.com/SYSTRAN/fuzzy-match. During the evaluation, there is no filter applied before scoring the segments with $\operatorname{Lev}$. However, the examples associated to the training samples are selected with $\operatorname{Lev}$and a BM25 filter. This ensures a reasonable time to build the set of examples for traning.
B.2 FAISS
The execution of the CLIR techniques is done with library FAISS [Douze et al., 2024]. The $k$NN search is performed on GPU (IndexFlatIP), without IVF (for faster search) or quantizer (for lower memory footprint) to ensure the retrieval of the best matches. Indeed, both optimizations can prevent the model from finding the optimal $k$ nearest neighbors.
Appendix C NMT Architectures Configuration
Both TM3-LevT and NFA models rely on the default backbone Transformer architecture [Vaswani et al., 2017]. We train one instance of TM3-LevT for each language pair (en-fr and de-en), using only the training data available in the corresponding datasets. Our NFA models are trained on much larger sets of parallel data, including a good share of Web data. We use these models to showcase the effectiveness of cross-lingual retrieval for RAMNT systems trained with large, high-quality, parallel datasets (8M sentences for en-fr and 11.5M for de-en). After inspecting the training datasets of NFA models, we found that some of them partly overlap with the training data; the corresponding contamination rates are reported in Table 4.
As we compare retrieval techniques for fixed MT architectures, these overlaps do not affect on our main conclusions (see also the discussion in §6.2).
In addition, we use the multilingual NLLB model [Costa-jussà et al., 2024] out-of-the-box for both language pairs.101010version: NLLB-200-distilled-1.3B
https://huggingface.co/facebook/nllb-200-1.3B
This large encoder-decoder model is used to generate the back-translation data; it also provides a baseline against which to appreciate the performance of RANMT architectures. As for the NFA models, the training data of NLLB partly overlaps with our test sets, for domains ECB, EMEA, JRC-Acquis and OpenSubtitles. Contamination is discussed in §6.2.
Appendix D Computational Cost of Retrieval
The computational cost can be dived into two categories: a fixed cost (independant from the number of sentences to translate) and a variable cost (proportional to this number). The retrieval process for training falls into the former, as well as the encoding of the whole set of monolingual target segments (for CLIR techniques) and the indexation of the TM (for fuzzy matching techniques). The variable costs lie in the fuzzy match search with filtering and the computation of $\operatorname{Lev}$, and in the CLIR search with the encoding of the source sentences, then the $k$NN search. In our experiments, we removed the filter (BM25) so that we obtain the best possible fuzzy matches (fuzzy-src, fuzzy-bt and fuzzy-gold), at the cost of a $\sim$100 times higher latency. However, when applying a BM25 filter to preselect 100 segments, we observe that both fuzzy matching and CLIR techniques can retrieve their closest match in about $\sim$1ms. We obtained this result un corpus ECB (150K sentences) with optimmal conditions:
  • the fuzzy matching is performed on a 8 cores CPU, thus parallelizing the search on 8 simultaneous source sentences;
  • as for CLIR, source sentences are encoded in batches of size 50, and the $k$NN search is handled by library FAISS. We used a V100-32GB GPU.

### v2

@@
-IJCAI–25 Formatting Instructions
+Improving Retrieval-Augmented Neural Machine Translation with Monolingual Data
 Abstract
 Conventional retrieval-augmented neural machine translation (RANMT) systems leverage bilingual corpora, e.g., translation memories (TMs).
-Yet, in many settings, in-domain monolingual target-side corpora are often available.
-This work explores ways to take advantage of such resources by retrieving relevant segments directly in the target language, based on a source-side query.
+Yet, in many settings, monolingual corpora in the target language are often available.
+This work explores ways to take advantage of such resources by directly retrieving relevant target language segments, based on a source-side query.
 For this, we design improved cross-lingual retrieval systems, trained with both sentence level and word-level matching objectives.
-In our experiments with two RANMT architectures, we first demonstrate the benefits of such cross-lingual objectives in a controlled setting, obtaining translation performances that surpass standard TM-based models.
-We then showcase our method on a real-world set-up, where the target monolingual resources far exceed the amount of parallel data and observe large improvements of our new techniques, which outperform both the baseline setting, and general-purpose cross-lingual retrievers.
+In our experiments with three RANMT architectures, we assess such cross-lingual objectives in a controlled setting, reaching performances that match those of standard TM-based models.
+We also showcase our method on a real-world settings, using much larger monolingual corpora, and observe strong improvements over both the baseline setting and general-purpose cross-lingual retrievers.
+Improving Retrieval-Augmented Neural Machine Translation with Monolingual Data
+Maxime Bouthors   Josep Crego   Dakun Zhang   François Yvon SYSTRAN by ChapsVision, 5 rue Feydeau, F-75002 Paris, France Sorbonne Université, CNRS, ISIR, F-75005 Paris, France {mbouthors, jcrego, dzhang}@chapsvision.com francois.yvon@cnrs.fr
 1 Introduction
@@
 Such techniques have been transposed in Statistical MT (Koehn and Senellart, 2010), more recently in Retrieval-Augmented Neural Machine Translation (RANMT) e.g., (Gu et al., 2018), making the entire process fully automatic.
-Most subsequent work has continued to leverage parallel data in the form of TMs, performing retrieval in the source language, then using the retrieved target side text in generation.
-Similar principles underlie few-shot MT with large language models, where both the source and target sides of examples are inserted into the prompt Brown et al.
-(2020).
+Most subsequent work has continued to leverage parallel data, performing retrieval in the source language, then using the retrieved target side text in generation.
+Similar principles underlie few-shot MT with large language models, where both the source and target sides of examples are inserted into the prompt (Brown et al., 2020).
 TM-based approaches typically retrieve examples with lexical fuzzy matchers, using BM25 and/or the Levenshtein distance (LED) as filters or rankers (Bouthors et al., 2024).
-These are known to be computationally efficient, and to often surpass continuous-space retrievers (Xu et al., 2020): this is because lexical (source) matches often come with appropriate translation of terms.
-However, relying on source-side lexical similarity raises two issues that can cause irrelevant segments to be retrieved: (a) noisy alignments between source and target side of a parallel instance; (b) intrinsic morphosyntactic divergences between the two languages (Dorr, 1994).
-These problems can be sidestep by performing retrieval directly on the target side, using Cross-Lingual Information Retrieval (CLIR) techniques.
-An important additional benefit would be to dispense with the use of parallel data and to potentially inform the generation process with relevant monolingual resources.
-As demonstrated by the recurrent use of target side monolingual data thanks to back-translation Sennrich et al.
-(2016), such resources are much easier to find than parallel ones in many domains, and also devoid of so-called translationese phenomena Bogoychev and Sennrich (2020).
-As such, monolingual corpus constitute a very useful, yet under-studied, source of high-quality segments for RANMT.
+These are known to be computationally efficient and to often surpass continuous-space retrievers (Xu et al., 2020): this is because lexical (source) matches often come with appropriate translation of terms.
+However, using source-side lexical similarity raises two issues that may cause inadequate segments to be retrieved: (a) noisy alignments between source and target side of a parallel instance; (b) intrinsic linguistic divergences between the two languages (Dorr, 1994).
+These problems can be sidestepped by performing retrieval directly on the target side, using Cross-Lingual Information Retrieval (CLIR) techniques (Cai et al., 2021; Tamura et al., 2023).
+An important additional benefit is to dispense with the use of parallel data and potentially inform the generation process with relevant monolingual resources.
+As demonstrated by the recurrent use of target side monolingual data thanks to back-translation (Sennrich et al., 2016a), such resources are much easier to find than parallel ones in many domains, and are also devoid of so-called translationese phenomena (Bogoychev and Sennrich, 2020).
+As such, monolingual corpora constitute a very useful, yet under-studied, source of high-quality segments for RANMT.
 To search such resources, cross-lingual retrievers are readily available thanks to multilingual sentence encoders such as LASER (Artetxe and Schwenk, 2019b) or LaBSE (Feng et al., 2022).
-They however lack the critical ability to ensure that the retrieved examples will have a large lexical overlap with the source text.
+As they only rely on dense representations, such retrievers however often fail to retrieve lexically relevant examples, i.e, examples that would contain translations of the exact source words.
 In this work, we propose novel methods to boost the lexical matching abilities of language-agnostic sentence encoders, thereby improving their effectiveness for RANMT.
 For this, we consider various ways to fine-tune a pre-trained encoder into reproducing the behaviour of lexical fuzzy matchers.
-We experiment with two language pairs (English-French and German-English) and multiple domains of various sizes and diversities.
-We consider two types of RANMT systems: an autoregressive source-augmented transformer (Bulte and Tezcan, 2019) and a non-autoregressive edit-based model (Bouthors et al., 2023).
+We experiment with three language pairs (English-French, German-English and English-Ukrainian) and multiple domains of various sizes and diversities.
+We consider three types of RANMT systems: an autoregressive source-augmented transformer (Bulte and Tezcan, 2019), a non-autoregressive edit-based model (Bouthors et al., 2023) and a Large Language Model (LLM), with in-context ($k$-shot) learning (Brown et al., 2020).
 We first study the performance of CLIR techniques in a controlled setting, showing that they match standard TM-based models with source-side retrieval.
-We then consider a much larger set-up, where the target monolingual resources far exceed the amount of parallel data and observe large improvements using our new techniques, which outperform both the baseline setting, as well as generic cross-lingual retrievers.
+We then consider much larger setups, where the target monolingual resources far exceed the amount of parallel data and observe large improvements using our new techniques, which outperform various baseline systems, as well as RANMT architectures relying on generic cross-lingual retrievers.
 2 Related Work
+Retrieval Augmented NMT
 Our work falls under the scope of Retrieval-Augmented Generation, leveraging both lexical (BM25 (Robertson and Walker, 1994), Levenshtein distance (Levenshtein, 1965)) and semantic methods for the Information Retrieval (IR) component.
 Semantic methods usually rely on dual encoders trained on parallel data (Gillick et al., 2018) with a contrastive loss (Sohn, 2016).
-Our focus is on massively multilingual sentence encoders that can encode mutual translations as close neighbors, such as e.g, LASER (Artetxe and Schwenk, 2019b) and LaBSE (Feng et al., 2022).
-In recent years, there has been a growing focus on integrating retrieval components into NMT systems.
+Our focus is on massively multilingual sentence encoders that encode mutual translations as close neighbors, such as LASER (Artetxe and Schwenk, 2019b) and LaBSE (Feng et al., 2022).
+In recent years, retrieval-augmented MT systems have received increasing attention.
 One key motivation is to enhance model transparency by providing users with the retrieved examples that inform the generated output (Rudin, 2019).
-A simple implementation of this idea encodes the target side of the example(s) along with the source sentence, providing an additional translation context (Bulte and Tezcan, 2019; Gu et al., 2018; Xia et al., 2019; He et al., 2021; Cheng et al., 2022; Agrawal et al., 2023).
-It is also possible to leverage both the source and target sides of the example(s) (Pham et al., 2020; Reheman et al., 2023).
-Instead of regenerating a complete text with an extended context, edit-based models aim to perform minimal changes to the retrived examples and generate their output in a non-autoregressive (NAR) fashion.
-Such techniques are studied in (Niwa et al., 2022; Xu et al., 2023; Zheng et al., 2023), adapting the Levenshtein transformer (Gu et al., 2019).
-This approach is extanded to handle multiple examples that are simultaneously edited in (Bouthors et al., 2023).
-MT systems based on Large Language Models (LLMs) can also seamlessly accomodate examples though In-Context Learning (ICL), where the LLM prompt is enriched with multiple demonstrations of the translation task, comprising both the source and target sides of parallel sentences (Radford et al., 2019).
-Several studies have focused on optimizing the performance of such architectures for NMT, examining the impact of prompt modifications, the number of demonstrations and the retrieval procedure (Moslem et al., 2023; Vilar et al., 2023; Zhang et al., 2023; Hendy et al., 2023; Bawden and Yvon, 2023; Zebaze et al., 2024).
-Our lexical alignment objective corresponds to a form of neural metric learning, where a high dimensional distance space is reduced to a smaller one (Suárez et al., 2021).
-This field of research aims to learn distances from the data (Kulis, 2013; Bellet et al., 2015), which is particulary useful for similarity-based algorithms used in IR (Cakir et al., 2019).
+A simple implementation of this idea encodes the target side of the example(s) along with the source sentence, providing an additional translation context (Gu et al., 2018; Bulte and Tezcan, 2019; Xia et al., 2019; He et al., 2021; Cheng et al., 2022; Agrawal et al., 2023).
+It is also possible to leverage both the source and target sides of the retrieved example(s) (Pham et al., 2020; Reheman et al., 2023).
+Rather than regenerating a complete translation with extended context, edit-based models perform minimal changes to the retrieved examples and generate their output in a non-autoregressive (NAR) fashion.
+Such models are studied in (Niwa et al., 2022; Xu et al., 2023; Zheng et al., 2023), adapting the Levenshtein Transformer (Gu et al., 2019).
+Bouthors et al.
+(2023) generalize this approach to simultaneously handle multiple examples.
+MT systems based on Large Language Models (LLMs) can also seamlessly accomodate examples through In-Context Learning (ICL), where the LLM prompt is enriched with demonstrations of the translation task, comprising both the source and target sides of parallel samples (Radford et al., 2019).
+Several studies have tried to optimize the performance of such architectures for MT, examining the impact of prompt modifications, the number of demonstrations and the retrieval procedure (Moslem et al., 2023; Vilar et al., 2023; Zhang et al., 2023; Hendy et al., 2023; Bawden and Yvon, 2023; Zebaze et al., 2024).
+Monolingual Data in NMT
+Augmenting NMT models with target side monolingual data by back-translating segments is a well-established strategy for improving model training (Sennrich et al., 2016a; Edunov et al., 2018), already considered in RANMT setup (Tezcan et al., 2024).
+(Cai et al., 2021; Tamura et al., 2023) dispense with back-translation, performing directly the retrieval in the target language with CLIR techniques: the latter uses existing retrievers, while the former jointly trains the retriever and the MT components.
+Metric Learning
+Our lexical alignment objectives (§ 3.2) corresponds to a form of neural metric learning, where a high-dimensional metric space is reduced to a smaller one (Suárez et al., 2021).
+This line of work aims to learn distances from the data (Kulis, 2013; Bellet et al., 2015), which is particulary useful for similarity-based algorithms used in IR (Cakir et al., 2019).
 3 Method
 3.1 Problem Formulation
-Given a source sentence $\mathbf{x}$ and a corpus of segments $\mathcal{D}=(\mathbf{y}_{1},\dots,\mathbf{y}_{N})$ in the target language, the goal of CLIR in NMT is to retrieve $k$ relevant examples: (~$(\tilde{}\mathbf{y}_{1},\dots,\tilde{}\mathbf{y}_{k})$.
+Given a source sentence $\mathbf{x}$ and a corpus of target segments $\mathcal{D}=(\mathbf{y}_{1},\dots,\mathbf{y}_{N})$, the goal of CLIR is to retrieve $k$ relevant examples: $(\tilde{}\mathbf{y}_{1},\dots,\tilde{}\mathbf{y}_{k})$.
 Our approach closely follows neural-based cross-lingual retrievers, and comprises two main steps.
 First, $\mathcal{D}$ is encoded with a multilingual sentence encoder $E_{\theta}$.
-Then, for any query source sentence $\mathbf{x}$, retrieval is performed as a $k$NN similarity search in the embedding space, using the cosine similarity:
+Then, for any query source sentence $\mathbf{x}$, retrieval is performed as a kkNN similarity search in the embedding space, using the cosine similarity:
 $\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y})=\cos(E_{\theta}(\mathbf{x}),E_{\theta}(\tilde{}\mathbf{y})).$ (1)
@@
 Usually, $E_{\theta}$ is trained as a Siamese encoder with a contrastive loss on parallel data, a usual IR objective (Sohn, 2016) that primarily captures semantic similarity.
-However, it is desirable for the retrieved $\tilde{\mathbf{y}}_{i}$ instances to exhibit surface-level similarities to what could potentially be a translation $\mathbf{y}$ of the source $\mathbf{x}$.
+However, it is desirable for the retrieved $\tilde{\mathbf{y}}_{i}$ instances to exhibit lexical similarities with the unknown translation $\mathbf{y}$ of the source $\mathbf{x}$.
 In specialized domains (like law or medicine), adherence to domain-specific terminology takes precedence over semantic similarity, thereby motivating the design of our proposed lexical-based alignment objective.
 To address this need, we study various fine-tuning objectives based on the Levenshtein similarity.
-These objectives are specifically designed to boost the retrieval of segments with a large numbers of lexical matches in the source query, by learning a new embedding space, achieved through the optimization of parameters $\theta$, such that:
+These objectives are specifically designed to boost the retrieval of segments with a large numbers of lexical matches in the source query, by learning a new embedding space, achieved through the optimization of parameters $\theta$, such that: [^1] As edit distances are only computed monolingually, our method does not assume any lexical overlap between $\mathbf{x}$ and $\tilde{}\mathbf{y}$.
 $f(\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}))\approx\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}),$ (2)
 where $f$ is a (potentially learnt) mapping from the codomain of the cosine function into the codomain of Levenshtein similarity function $\operatorname{Lev}$, defined as:
-$\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})=1-\frac{\Delta(\mathbf{y},\tilde{}\mathbf{y})}{\max(|\mathbf{y}|,|\tilde{}\mathbf{y}|)},$ (3)
-where $\Delta$ is the Levenshtein distance (Levenshtein, 1965).
+$\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})=1-\frac{\Delta(\mathbf{y},\tilde{}\mathbf{y})}{\max(|\mathbf{y}|,|\tilde{}\mathbf{y}|)}.$ (3)
+$\Delta$ is the Levenshtein distance (Levenshtein, 1965).
 3.2 Training
-We consider three objective functions. The first two learn to predict the Levenshtein similarity. We choose $f$ to be a learnable fonction that maps $[-1,1]$ to $[0,1]$:
+We consider three objective functions. The first two learn to predict the Levenshtein similarity (3). We choose $f$ to be a learnable function that maps [1,1][-1,1] to $[0,1]$:
 $f(t)=\sigma(a\times\operatorname{arctanh}(t)+b),$ (4)
 with $a$ (slope) and $b$ (position) parameters and $\sigma$ the sigmoid function. The loss is defined as:
-$\mathcal{L}(\mathbf{x},\mathbf{y},\tilde{}\mathbf{y})=\operatorname{Err}(f(\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y})),\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})),$ (5)
-where the error function $\operatorname{Err}$ is either the mean square error (MSE) or the mean absolute error (MAE) defined as:
+$\hskip-5.0pt\mathcal{L}(\mathbf{x},\mathbf{y},\tilde{}\mathbf{y})=\operatorname{Err}(f(\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y})),\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})),$ (5)
+where the error function $\operatorname{Err}$ is either the mean square (MSE) or the mean absolute error (MAE):
 $\operatorname{MSE}(x,y)=(x-y)^{2}$ (6)
 $\operatorname{MAE}(x,y)=|x-y|.$ (7)
-The third objective is a “learning to rank” objective (Cao et al., 2007).
-Given a set of $k$ retrieved segments ~$\tilde{}\mathbf{y}_{[1:k]}$ sorted in decreasing order w.r.t.
+The third objective is a “learning to rank" objective (Cao et al., 2007).
+Given a set of $k$ retrieved segments $\tilde{}\mathbf{y}_{[1:k]}$ sorted in decreasing order w.r.t.
 $\operatorname{Lev}(\mathbf{y},\cdot)$, it computes:
-$\mathcal{L}(\mathbf{x},\mathbf{y},\tilde{}\mathbf{y}_{[1:k]})=$ $\sum_{i>j}\max(0,\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{i})-\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{j})+$
-$m\times|\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})-\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})|)$ (8)
+$\mathcal{L}($ $\mathbf{x},\mathbf{y},\tilde{}\mathbf{y}_{[1:k]})=\sum_{i>j}\max(0,U_{ij})$ (8)
+$U_{ij}$ $=\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{i})-\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{j})$
+$+m\times|\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})-\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})|)$ (9)
 with $m$ a scaling factor.
-Equation (8) defines a pairwise margin loss, [^1] Derived from the triplet loss (Schultz and Joachims, 2003).
+Equation (9) defines a pairwise margin loss, [^2] Inspired by the triplet loss (Schultz and Joachims, 2003).
 with an adaptive margin proportional to the absolute difference between the two individual $\operatorname{Lev}$ similarities.
@@
 4.1 Data
-We consider two translation tasks: from English to French (en-fr) and from German to English (de-en).
-Each test corpus contains a wide variety of textual domains (16), all from the Opus corpus (Tiedemann, 2012).
-The details are shown in appendix A.
-We also prepare a large corpus of Wikipedia pages in French.
+We consider three translation directions: English into French (en-fr), German into English (de-en) and English into Ukrainian (en-uk).
+For our controlled experiments, we consider a wide variety of textual domains (16), using public parallel datasets from the Opus corpus (Tiedemann, 2012).
+Details are in appendix A.
+For the large scale experiments, we expand the English-French setting with a large corpus of Wikipedia pages in French.
 [^3] https://huggingface.co/datasets/Plim/fr_wikipedia_processed.
 We split the text in pseudo-sentences via regular expression matching and remove duplicates as well as fuzzy matches from the valid/test sets [^4] Each best match with $\operatorname{Lev}>0.9$ is simply removed.
-of the parallel Wikipedia corpus split, ending up with about $45$M French segments.
+of the parallel Wikipedia corpus, ending up with about 4545M French segments.
+The corresponding English-Ukrainian experiments, reported in Appendix H, reproduce the setting of Tezcan et al.
+(2024), using texts in the legal domain.
 4.2 Metrics
@@
 (3)) between the 1-best match and the reference, for sentences in the validation set, with retrieval performed from the entire train set.
-We additionally report xsim error rates, which measure the abililty to identify matched sentences in a bilingual parallel corpus (Artetxe and Schwenk, 2019a).
-We assess translation quality with BLEU scores (Papineni et al., 2002) computed with SacreBLEU (Post, 2018),444signature: nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.1.0; and COMET-20555Unbabel/wmt22-comet-da, using the official implementation.
-scores (Rei et al., 2020).
+We also report xsim error rates, which measure the abililty to identify matched sentences in a bilingual parallel corpus (Artetxe and Schwenk, 2019a).
+We assess translation quality with BLEU (Papineni et al., 2002) computed with SacreBLEU (Post, 2018),555signature: nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.1.0; as well as COMET-20666Unbabel/wmt20-comet-da, using the defaults settings.
+(Rei et al., 2020).
 5 Experimental Settings
 5.1 Retrieval Settings
-Our controlled experiments use parallel corpora in two language pairs (en-fr and de-en).
-This allows us to perform a head-to-head comparison between standard source-side fuzzy matching and our novel cross-lingual retrieval methods.
-In each configuration, we retrieve up to $k=3$ closest examples.
-We compare the following retrieval methods:
-  • fuzzy-src: source-to-source fuzzy matching ranked by their $\operatorname{Lev}$ similarity, as in standard RANMT settings;
+Our experiments consider the following retrieval methods, with up to $k=3$ closest examples:
+  • fuzzy-src: source-to-source fuzzy matching ranked by their $\operatorname{Lev}$ similarity, as in baseline RANMT settings;
   • fuzzy-gold: target-to-target fuzzy matching, using the same procedure as fuzzy-src.
-This setup requires access to reference target translations and corresponds to an oracle score, giving an empirical upper-bound of what an ideal retriever could achieve;
-  • fuzzy-bt: the target side of training segments are back-translated into the source language Sennrich et al.
-(2016), then used for retrieval as in fuzzy-src.
-This simulates an alternative to cross-lingual retrieval in situations where only target side examples are available;
+This retriever requires access to reference target translations and corresponds to an oracle setting, giving an empirical upper-bound of what an ideal system could achieve.
+It thus is only reported for the controlled experiments of § 6.1 and § 6.2;
+  • fuzzy-bt: the target side of training segments are back-translated using NLLB (Costa-jussà et al., 2024) [^7] We used NLLB-200-distilled-1.3B out-of-the-box.
+    https://huggingface.co/facebook/nllb-200-1.3B
+    into the source language Sennrich et al.
+(2016a), then used for retrieval as in fuzzy-src.
+This simulates an alternative to cross-lingual retrieval in situations where only target side examples are available; [^8] Back-translation is further discussed in § G.
   • dense: a BERT base Devlin et al.
-(2019) model trained from scratch with a contrastive loss (Sohn, 2016), usable for source-to-target retrieval.
-Given a batch $\mathcal{B}=((\mathbf{x}_{1},\mathbf{y}_{1}),\dots,(\mathbf{x}_{n},\mathbf{y}_{n}))$, the loss is computed as:
-    $\mathcal{L}(\mathcal{B})=-\sum_{i}\log\frac{\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{i}))}{\sum_{j}\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{j}))};$ (9)
+(2019) model for source-to-target retrieval.
+It is trained from scratch with a contrastive loss (Sohn, 2016).
+The loss $\mathcal{L}(\mathcal{B})$ for batch $\mathcal{B}=((\mathbf{x}_{1},\mathbf{y}_{1}),\dots,(\mathbf{x}_{n},\mathbf{y}_{n}))$ is defined as:
+    $\sum_{i}\log\frac{\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{i}))}{\sum_{j}\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{j}))};$ (10)
   • dense+bow: we add a bag-of-word loss to dense as in (Cai et al., 2021) to enforce lexical information in cross-lingual encodings.
-Given a pair of parallel sentences $(\mathbf{x},\mathbf{y})$, the loss is computed as:
-    $\mathcal{L}(\mathbf{x},\mathbf{y})=-\sum_{w\in\mathcal{Y}}\log p_{\theta_{1}}(w|\mathbf{x})-\sum_{w\in\mathcal{X}}\log p_{\theta_{2}}(w|\mathbf{y}),$ (10)
-    where $\mathcal{X}$, $\mathcal{Y}$ are the bag-of-words of $\mathbf{x}$ et $\mathbf{y}$; $p_{\theta_{1}}$ and $p_{\theta_{2}}$ are distinct linear projections of $E_{\theta}(\mathbf{x})$ and $E_{\theta}(\mathbf{y})$;
+Given a pair of parallel sentences $(\mathbf{x},\mathbf{y})$, the loss $\mathcal{L}(\mathbf{x},\mathbf{y})$ is computed as:
+    $\sum_{w\in\mathcal{Y}}\log p_{\theta_{1}}(w|\mathbf{x})-\sum_{w\in\mathcal{X}}\log p_{\theta_{2}}(w|\mathbf{y}),$ (11)
+    where $\mathcal{X}$, $\mathcal{Y}$ are the bag-of-words for $\mathbf{x}$ et $\mathbf{y}$; $p_{\theta_{1}}$ and $p_{\theta_{2}}$ are distinct linear projections of $E_{\theta}(\mathbf{x})$ and $E_{\theta}(\mathbf{y})$;
   • LASER: we use LASER (Artetxe and Schwenk, 2019b) as a first pre-trained source-to-target retriever.
 LASER derives from a multilingual RNN-based translation model, repurposing the source-side embeddings to define a similarity measure between multilingual sentence representations;
-  • LaBSE: Similar to dense, LaBSE (Feng et al., 2022) uses a dual encoder model based on the BERT base architecture and is trained with a contrastive loss; it additionnally benefits from pre-training on large sets of monolingual data, and covers more than 110 languages;
-[…]
-In our setting, fine-tuning updates all the model parameters.
-The global learning rate is 1e-4, as for $a$ and $b$ (eq.
+  • LaBSE: Similar to dense, LaBSE (Feng et al., 2022) uses a dual encoder model based on the BERT base architecture and is trained with a contrastive loss; it additionally benefits from pre-training on large sets of monolingual data, and covers more than 110 languages;
+  • ft-[model]-[Err]: we fine-tune either one of the aforementioned encoders with the losses of §3.2; Err is one of $\{\text{MSE},\text{MAE},\text{Rank}\}$ (e.g. ft-LaBSE-MSE).
+In our setting, fine-tuning updates all the parameters.
+The global learning rate is 1e-4; as for $a$ and $b$ (eq.
 4) it is set to 1e-2.
-Each source $\mathbf{x}$ is presented with three ~$\tilde{}\mathbf{y}$ (eq.
+Each source $\mathbf{x}$ is presented with three $\tilde{}\mathbf{y}$ (eq.
 5) determined by the top 3 scoring matches obtained with dense+bow.
 The validation score is an in-batch normalized discounted cumulative gain (NDCG) score (Wang et al., 2013).
-The retrieval is always performed in an in-domain setting.
+Retrieval is always performed in an in-domain setting, ensuring that translations for a specific domain are exclusively based on examples from that domain.
 For all methods, a threshold is further applied to filter out low similarity segments.
-For each retrieval method, we adjust this threshold on the validation set to achieve an averaged retrieval rate [^6] Proportion of sentences with at least one retrieved example.
-of $50$%.
-Using a constant retrieval rate for all retrievers ensures that no system is unduly rewarded for retrieving a greater number of similar examples, and that our comparisons of retrievers only reflects differences in the quality of the TM matches.
-Please, refer to appendix B for the complete search configurations.
+For each retrieval method, we adjust this threshold on the validation set to achieve an averaged retrieval rate [^9] Fraction of sentences with at least one retrieved example.
+of 5050%.
+Using a constant retrieval rate for all retrievers ensures that no system is unduly rewarded for retrieving a greater number of similar examples, so that our comparisons of retrievers only reflect differences in the quality of the TM matches.
+Refer to appendix B for a discussion of search configurations.
 5.2 NMT Architectures
-We consider two RANMT architectures.
+We consider three RANMT architectures.
 The first is an in-house implementation of NFA (Bulte and Tezcan, 2019), which extends a basic encoder-decoder MT architecture by prepending the target side of TM examples to the source text on the encoder side.
@@
 Retrieved examples used in training are also selected based on source-side similarity measures; we use up to $k=3$ examples.
+The third architecture is EuroLLM-9B (Martins et al., 2025), specifically designed and trained for multilingual tasks such as Machine Translation.
+We fine-tune the EuroLLM-9B base model on the en-fr and de-en training corpora using a prompt that only contains relevant examples in the target language.
+A discussion of the adaption of EuroLLM for this setup is in Appendix D.
 6 Results
-We first report the results of controlled experiments, in order to contrast mono-lingual vs.
-cross-lingual retrieval strategies.
-We then analyze the results of a larger-scale setting, where we can only exploit monolingual data.
+We first run controlled experiments serving two main purposes: (a) to compare the quality of various retrieval models (§ 6.1); (b) to evaluate the gap between the baseline source-side retrieval and our proposed CLIR-based approaches (§ 6.2), an analysis that crucially requires parallel datasets.
+These experiments use two language pairs (en-fr and de-en).
+This configuration is somewhat ideal, as CLIR methods are primarily meant for cases where only monolingual data are available, as illustrated by the experiments in § 6.3.
 test en-fr test de-en
-$\operatorname{Lev}$uparrow xsim $\downarrow$ $\operatorname{Lev}$uparrow xsim $\downarrow$
+$\operatorname{Lev}$ $\uparrow$ xsim $\downarrow$ $\operatorname{Lev}$ $\uparrow$ xsim $\downarrow$
 fuzzy-gold (oracle) 50.0 - 53.4 -
@@
 ft-LaBSE-Rank 36.0 7.3 34.2 23.3
-Table 1: Average retrieval scores (xsim error rate and average Levenshtein similarity), averaged for all domains.
+Table 1: Average retrieval scores (xsim error and Levenshtein similarity), averaged accross domains.
 6.1 Retrieval Scores
@@
 For en-fr, source-side fuzzy matching yields more similar segments than CLIR techniques, a gap that vanishes after fine-tuning with MAE for dense+bow and even more so for LaBSE.
-For de-en, the source-side fuzzy matching baseline is much worse, revealing a large gap with the oracle condition, which might be due to residual misalignments or to morphological differences between German and English.
-[^7] Inflection phenomena in German create multiple variants for each lexeme, making exact matches at the word level much less likely than in English.
+For de-en, the gap between the source-side fuzzy matching baseline and the oracle condition is much larger than for en-fr, which might be due to residual misalignments or to morphological differences between German and English.
+[^10] Inflectional processes in German create multiple variants for each lexeme, making exact matches at the word level much less likely in German than in English.
 All CLIR techniques widely outperform the monolingual settings (fuzzy-src and fuzzy-bt).
-Here, however we do not see much gain from fine-tuning, perhaps owing to a smaller fune-tuning dataset.
-As expected, fine-tuning with a lexical loss is slightly detrimental to the sentence-level retrieval scores (xsim error) for both dense+bow and LaBSE.
+For this language pair, however we do not see much gain from fine-tuning, perhaps owing to a smaller fune-tuning dataset.
+As expected, fine-tuning with a lexical loss is always slightly detrimental to the sentence-level retrieval scores (xsim error) for both dense+bow and LaBSE.
 Training dense+bow along with a bag-of-word loss (eq.
@@
 Retrieval-free Baselines
-We use both NLLB and NFA without any retrieved examples as our machine translation baselines.
+We use NLLB, NFA and EuroLLM without any retrieved examples as machine translation baselines.
 Their corresponding BLEU scores, averaged across domains, are respectively 39.0 (en-fr) and 33.8 (de-en) for NLLB and 48.0 (en-fr) and 38.2 (de-en) for NFA.
 Details per domain are in Table 4.
-TM3-LevT NFA
-test en-fr test de-en test en-fr test de-en
-BLEU COMET BLEU COMET BLEU COMET BLEU COMET
-sans exemple - - - - 48.0 87.1 38.2 79.8
-fuzzy-gold (oracle) 45,3 51,7 31,3 -8,6 52,2 87,8 44,7 82,2
-fuzzy-src 43,8 48,4 28,8 -10,3 51,3 87,5 41,9 81,7
-fuzzy-bt 40,5 43,8 22,0 -22,9 49,9 87,2 38,9 81,3
-dense 41,5 42,2 - - 49,9 87,1 - -
-dense+bow 42,4 44,1 28,3 -11,6 50,5 87,3 40,0 80,9
-ft-dense+bow-MSE 43,1 46,0 29,0 -12,7 51,0 87,5 41,6 81,5
-ft-dense+bow-MAE 43,2 46,3 29,1 -12,6 51,0 87,4 41,8 81,4
-ft-dense+bow-Rank 42,1 44,3 26,0 -17,9 50,7 87,3 41,8 80,9
-LASER 42,5 44,3 30,4 -7,7 49,9 87,3 41,1 81,3
-LaBSE 43,1 45,5 29,8 -9,5 50,5 87,3 41,5 81,4
-ft-LaBSE-MSE 43,7 47,8 27,7 -12,5 51,2 87,5 41,6 81,6
-ft-LaBSE-MAE 43,6 47,7 29,1 -9,2 51,2 87,5 42,0 81,5
-ft-LaBSE-Rank 43,3 47,1 28,4 -12,2 51,1 87,5 41,8 81,6
-Table 2: Average translation scores: TM3-LevT and NFA, all domains.
+TM3-LevT NFA EuroLLM
+test en-fr test de-en test en-fr test de-en test en-fr test de-en
+BLEU COMET BLEU COMET BLEU COMET BLEU COMET BLEU COMET BLEU COMET
+no example - - - - 48.0 87.1 38.2 79.8 44.4 86.3 40.7 82.6
+fuzzy-gold (oracle) 45.3 51.7 31.3 -8.6 52.2 87.8 44.7 82.2 52.6 88.1 46.8 83.9
+fuzzy-src 43.8 48.4 28.8 -10.3 51.3 87.5 41.9 81.7 51.6 87.8 41.6 83.0
+fuzzy-bt 40.5 43.8 22.0 -22.9 49.9 87.2 38.9 81.3 46.4 86.9 41.8 82.9
+dense 41.5 42.2 - - 49.9 87.1 - - 50.4 87.4 - -
+dense+bow 42.4 44.1 28.3 -11.6 50.5 87.3 40.0 80.9 51.1 87.5 44.1 83.0
+ft-dense+bow-MSE 43.1 46.0 29.0 -12.7 51.0 87.5 41.6 81.5 51.3 87.8 45.1 83.3
+ft-dense+bow-MAE 43.2 46.3 29.1 -12.6 51.0 87.4 41.8 81.4 51.5 87.7 45.2 83.4
+ft-dense+bow-Rank 42.1 44.3 26.0 -17.9 50.7 87.3 41.8 80.9 50.4 87.5 43.6 83.1
+LASER 42.5 44.3 30.4 -7.7 49.9 87.3 41.1 81.3 50.9 87.4 44.9 83.1
+LaBSE 43.1 45.5 29.8 -9.5 50.5 87.3 41.5 81.4 51.5 87.6 45.1 83.2
+ft-LaBSE-MSE 43.7 47.8 27.7 -12.5 51.2 87.5 41.6 81.6 51.6 87.8 44.2 83.2
+ft-LaBSE-MAE 43.6 47.7 29.1 -9.2 51.2 87.5 42.0 81.5 51.5 87.8 44.9 83.3
+ft-LaBSE-Rank 43.3 47.1 28.4 -12.2 51.1 87.5 41.8 81.6 51.3 87.7 44.8 83.2
+Table 2: Average translation scores: TM3-LevT, NFA, EuroLLM, all domains.
+Significant results (p=0.05) w.r.t.
+LaBSE are marked with .
+Best CLIR results are in bold.
+When fuzzy-src is not outperformed, its score is in bold.
 English-French German-English
@@
 ft-LaBSE-Rank 64.1 62.6 39.8 58.1 66.3 45.3 50.0 47.5 44.0 46.0 38.1 43.7 15.9 61.0 57.2 31.2
-Table 4: Per domain BLEU scores for NFA systems and NLLB.
+Table 4: Per domain BLEU scores for NLLB and NFA systems.
+Contaminated domains for NLLB (used in back-translation) are underlined.
 The contamination rate (top line) for NFA is the percentage of overlap of its training data with the test set.
-Contaminated domains for NLLB (used in back-translation) are underlined.
+English-French German-English
+ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
+no example 49.9 47.7 39.2 45.7 55.5 40.9 34.0 48.1 42.0 41.3 44.2 41.6 23.3 54.4 52.9 31.1
+fuzzy-gold 63.7 68.4 39.5 65.2 66.2 51.1 33.8 52.1 41.9 48.9 48.2 50.4 25.6 63.6 62.2 32.2
+fuzzy-src 62.6 66.9 39.5 63.3 64.9 49.0 33.7 51.6 41.7 47.4 47.0 43.3 23.6 55.1 54.3 31.8
+fuzzy-bt 52.0 54.1 39.3 49.1 56.5 43.2 33.8 50.4 41.5 45.0 46.1 42.5 23.6 54.8 56.4 31.9
+LASER 63.2 65.3 39.4 62.1 65.5 46.1 33.8 51.0 41.7 45.8 46.4 46.6 24.4 62.4 59.9 31.4
+LaBSE 63.5 66.1 39.4 63.3 65.8 48.4 33.8 51.2 41.6 46.4 46.5 47.2 23.7 62.7 60.3 31.4
+ft-LaBSE-MSE 62.8 66.3 39.4 63.2 64.9 50.0 33.8 51.7 41.7 46.8 47.3 46.6 24.1 60.0 58.7 31.5
+ft-LaBSE-MAE 62.8 65.9 39.4 63.3 65.1 49.6 33.8 51.5 41.7 46.8 47.1 47.6 24.2 61.3 59.9 31.6
+ft-LaBSE-Rank 62.4 65.3 39.4 62.7 64.6 49.3 33.8 51.5 41.7 47.3 46.7 47.5 23.8 61.1 60.2 31.5
+Table 5: Per domain BLEU scores for EuroLLM.
 TM3-LevT
-Translation results achieved with TM3-LevT are in Table 2.
+Translation results for TM3-LevT are in Table 2 (left part).
 The difference between the oracle (fuzzy-gold) and the baseline (fuzzy-src) gives an idea of the quality of the monolingual matches: the quality is pretty high in en-fr, owing to the cleaning procedure used for these data, more noisy in de-en, with a gap of about $2.5$ BLEU points.
@@
 NFA
-The corresponding results for NFA are in Table 2.
-This model is a much stronger baseline than TM3-LevT and the margin of progress (up to fuzzy-gold) is dimmer.
-Yet, we see that (a) retrieving examples directly on the target side (with dense, LASER, and LaBSE) already improves the example-free baseline by about $3$ BLEU points; (b) an extra step of fine-tuning of cross-lingual retrievers (with MSE and MAE) enables us to fully close the gap with source-side matching for both metrics and language pairs.
-In comparison, back-translating the monolingual target data also improves our baseline systems, but by a much smaller margin.
+The corresponding results for NFA are in Table 2 (middle part).
+This model is a much stronger baseline than TM3-LevT and the progress margins (with respect to fuzzy-gold) is dimmer.
+Yet, we see that (a) retrieving examples directly on the target side (with dense, LASER, and LaBSE) improves the example-free baseline by about $3$ BLEU points; (b) further fine-tuning the cross-lingual retrievers (with MSE and MAE) fully closes the gap with source-side matching for both metrics and language pairs.
+We observe that back-translating the monolingual target data also improves baseline systems, albeit by a much smaller margin.
+EuroLLM
+Results for the fine-tuned version of EuroLLM are in the right part of Table 2.
+This system turns out to be the most effective system overall, slightly outperforming NFA with $k=3$ TM examples.
+A first observation is that this model is particularly strong for translating into English (from German) where using a CLIR outperforms all baselines.
+For this model, the averaged gains achieved by fine-tuning the cross-lingual retrievers are very small, and vary accross domains (details in Table 5).
+For this system, the gap between the LaBSE system and the oracle condition (fuzzy-gold) is already small in the two language pairs, showcasing the very strong built-in cross-lingual abilities of this model; yet, the difference that remains suggests that lexical similarity should still matter.
 Per-domain analysis
-Per-domain BLEU scores are in Tables 3 (TM3-LevT) and 4 (NFA).
+Per-domain BLEU scores are in Tables 3 (TM3-LevT), 4 (NFA), and 5 (EuroLLM).
 The domains for which we expect CLIR methods to work well need to satisfy two conditions: (a) a large enough set of examples to retrieve from, so that the choice the retriever actually makes a difference; (b) a small diversity of texts in that domain, so that retrieved instances are sufficiently similar to the input.
-This is well reflected in Table 3 (TM3-LevT) where CLIR techniques (with fine-tuning in en-fr, without in de-en) match the performance of the baseline in all domains, and even outpeform it in specialized domains such as ECB (finance), JRC (law), KDE (information technology) and Wikipedia.
+This is well reflected for instance in Table 3 (TM3-LevT) where CLIR techniques (with fine-tuning in en-fr, without in de-en) match the performance of the baseline in all domains, and even outpeform it in specialized domains such as ECB (finance), JRC (law), KDE (information technology) and Wikipedia.
 We do not see the same gains in Ubuntu (very small TM) or Europarl (large TM, with a diverse set of texts).
 For NFA, we also observe that using CLIR techniques yields scores that match the baseline; for both monolingual or cross-lingual retrieval, the improvement margin, as defined by the difference with fuzzy-gold, is small for all domains, except for Koran (de-en), where CLIR techniques once again demonstrate their effectiveness.
+Regarding EuroLLM, we note that performing CLIR with the default version of LaBSE already yields near optimal performance for many domains.
+Fine-tuning has here a reduced effect on BLEU scores, sometimes slightly improving (e.g., for en-fr: EMEA, KDE, PHP, Wiki, UBU) or, more rarely, decreasing (e.g., for en-fr, ECB, JRC) performance.
+These variations highlight the fact that optimal translation scores ultimately also depend on the retrieval pool and the associated example quality.
 Comparison of fine-tuning losses
-The standard contrastive loss of equation (9), when applied in MT, primarily seeks to identify parallel sentences within collections of monolingual segments.
+The standard contrastive loss of eq.
+(10), when applied in MT, primarily seeks to identify parallel sentences within collections of monolingual segments.
 Yet, RANMT requires examples that are lexically similar to the source sentences.
-The bag-of-word loss of equation (10) effectively increases the rate of lexically similar sentences, thus enhancing the translation scores.
+The bag-of-word loss of eq.
+(11) effectively increases the rate of lexically similar sentences, thus enhancing the translation scores.
 As for our three fine-tuning strategies, MSE and MAE (eq.
@@
 On data contamination
-We identified two sources of data contamination.
+We identify two sources of data contamination.
 First, NFA training data partly overlaps with some of our test sets, in varying proportion (from 0 to 100) across domains (see Table 4).
@@
 In the per-domain results of Table 4, this sensitivity is assessed by the gap between example-free and fuzzy-gold scores.
-As this gap is large for most domains (with the exception of Europarl and News in en-fr, and Sub in de-en), we can still clearly assess the benefits of cross-lingual retrieval for the large majority of domains.
+As this gap is large for most domains (except Europarl and News in en-fr, and Sub in de-en), we can still clearly assess the benefits of cross-lingual retrieval for the large majority of domains.
 Second, the training data of NLLB, used for back-translation, also partly overlaps with our test, which may boost the efficiency of the fuzzy-bt baseline.
 Even for these domains (e.g., ECB, EMEA, JRC-Acquis), cross-lingual methods still achieve much higher BLEU scores than back-translation.
-Figure 1: Evolution of BLEU and COMET scores of TM3-LevT (top) and NFA (bottom) when growing the retrieval pool.
-6.3 A Large-scale Experiment
+Figure 1: BLEU and COMET scores with a growing retrieval pool: TM3-LevT (top), NFA (middle) and EuroLLM (bottom).
+6.3 Large-scale Experiments
 Retrieving in a monolingual corpus
-We use the models trained on the en-fr dataset to retrieve segments from the large monolingual Wikipedia dataset (see §4.1), using the Wikipedia test already used in the controled experiments set to assess performance.
-We use the same filtering threshold as for the controlled experiments to keep the settings comparable and ensure that the only difference is the increased size of the retriveal set.
+We use the models trained on the en-fr [^11] A comparable experiment, targeting translation from English into Ukrainian is in Appendix H; these results also highlight the benefits of CLIR over monolingual retrieval in parallel data, especially when using a fine-tuned version of LaBSE.
+dataset to retrieve segments from the large monolingual Wikipedia dataset (see §4.1), using the Wikipedia test already used in the controlled experiments set to assess performance.
+We use the same filtering threshold as for the controlled experiments to keep the settings comparable and ensure that the only difference is the increased size of the retrieval set.
 We only consider the variants of LaBSE, which achieve the best results in the small data condition.
@@
 The first observation is that increasing the retrieval set (from 200K segments in the controlled experiments – Tables 3 and 4 – to 45M segments) yields significant performance boosts across the board (e.g, +1.6 BLEU for NFA and fuzzy-bt, +0.4 BLEU points for NFA and LaBSE).
-With an additional lexical fine-tuning step, the benefits of the cross-lingual approach are again very clear, with large BLEU differences relative to the small data condition:+3.9 for ft-LaBSE-MAE with TM3-LevT, +3.4 when using NFA.
-TM3-LevT NFA
-BLEU COMET BLEU COMET
-no example - - 36.4 84.8
-fuzzy-bt 35.6 76.7 38.6 84.8
-LaBSE 37.7 77.2 40.1 85.0
-ft-LaBSE-MSE 38.3 77.2 40.9 84.8
-ft-LaBSE-MAE 38.7 77.6 41.2 84.9
-ft-LaBSE-Rank 38.5 77.5 40.8 85.0
-Table 5: BLEU and COMET scores with systems TM3-LevT and NFA with the Wikipedia retrieval pool.
+With an additional lexical fine-tuning step, the benefits of the cross-lingual approach are again very clear, with large BLEU differences relative to the small data condition:+3.9 for ft-LaBSE-MAE with TM3-LevT, +3.4 when using NFA, +2 using EuroLLM.
+TM3-LevT NFA EuroLLM
+BLEU COMET BLEU COMET BLEU COMET
+no example - - 36.4 84.8 44.2 86.3
+fuzzy-bt 35.6 76.7 38.6 84.8 47.1 86.4
+LaBSE 37.7 77.2 40.1 85.0 48.8 86.5
+ft-LaBSE-MSE 38.3 77.2 40.9 84.8 49.0 86.3
+ft-LaBSE-MAE 38.7 77.6 41.2 84.9 49.2 86.3
+ft-LaBSE-Rank 38.5 77.5 40.8 85.0 48.6 86.3
+Table 6: BLEU and COMET scores with systems TM3-LevT, NFA and EuroLLM with the Wikipedia retrieval pool.
 Effect of the retrieval pool size
-To better analyse the effect of the retrieval pool size, we progressively increase the Wikipedia retrieval set by batches of $500$K sentences from $500$K to $45$M, and compute the corresponding translation scores for each of the retrieval systems.
-Figure 1 reports the evolution of the translation metrics for LaBSE, its fine-tuned variants and fuzzy-bt.
+To better analyse the effect of the retrieval pool size, we progressively increase the Wikipedia retrieval set by batches of 500500K sentences from 500500K to 4545M, and compute the corresponding translation scores for each retriever.
+Figure 1 displays the evolution of the translation metrics for LaBSE, its fine-tuned variants and fuzzy-bt.
 Overall, for all models, the benefit of growing the retrieval pool yields an almost linear BLEU increase.
-COMET variations are much smaller, especially for NFA where it remains nearly constant for our best setups.
-For both metrics, ft-LaBSE-MAE outperforms the other setups for all data sizes.
+COMET variations are much smaller, especially for NFA and EuroLLM where it remains nearly constant.
+For both metrics, ft-LaBSE-MAE outperforms the other setups for all data sizes in most cases.
 fuzzy-bt consistently lags behind CLIR alternatives, especially with TM3-LevT.
 We also monitor in Table 7 two additional scores: the retrieval rate (percentage of segments for which at least one example is found) and the average $\operatorname{Lev}$ similarity between the reference and its closest retrieved neighbor.
-For all retrieval methods, we see large increases, meaning a larger number of retrieved examples is retrieved, and that their average quality (i.e their similarity to the reference) also improves, highliting the usefulness of monolingual data for this domain.
+For all retrieval methods, we see large increases, meaning a larger number of examples is retrieved and that their average quality (i.e, their similarity to the reference) also improves, highliting the usefulness of monolingual data for this domain.
 TM mono
-RR $\uparrow$ $\operatorname{Lev}$uparrow RR $\uparrow$ $\operatorname{Lev}$uparrow
+$R$ $\uparrow$ $\operatorname{Lev}$ $\uparrow$ $R$ $\uparrow$ $\operatorname{Lev}$ $\uparrow$
 fuzzy-bt 39.9 24.0 60.2 31.0
@@
 ft-LaBSE-Rank 25.7 26.7 61.2 36.7
-Table 6: Comparison of the retrieval rates (RR) and average Levenshtein similarity ($\operatorname{Lev}$) on the Wikipedia test set of the systems retrieving in the TM vs.
+Table 7: Comparison of the retrieval rates ($R$) and average Levenshtein similarity ($\operatorname{Lev}$) on the Wikipedia test set of the systems retrieving in the TM vs.
 in the large monolingual corpus (mono).
+The translation examples in Appendix F illustrate the benefits of enlarging the retrieval pools, notably because they provide longer lexical matches, thereby improving the generated text.
 7 Conclusions and Outlook
 In this paper, we have explored various approaches to take advantage of monolingual examples in retrieval-augmented neural machine translation.
-Our main conclusion, resulting from experiments on 2 language directions and 16 varied textual domains, is that retrieving directly in the target language can be as effective as fuzzy-matching on the source-side, especially with lexical-based fine-tuning objectives.
+Our main conclusion, resulting from experiments on 3 language directions and 16 varied textual domains, is that retrieving directly in the target language can be as effective as fuzzy-matching on the source-side, especially with lexical-based fine-tuning objectives.
 Using these techniques, we were also able to obtain large BLEU increases over strong baselines on the Wikipedia dataset (up to +3.8 BLEU points for our best configuration), showcasing the benefits of RANMT in a large-scale setup.
@@
 We identify several ways to improve these results: (a) by fine-tuning the retrieval models on very large multilingual datasets, adopting data settings that are comparable to the pre-training of LASER/LaBSE; (b) using more recent sentence embedders such as SONAR (Duquenne et al., 2023); (c) using CLIR techniques to train the RANMT models, instead of relying on source-side matches as we have done here; (d) revisiting the other aspect of the retrieval pipeline – e.g., relaxing the filtering threshold during example selection.
-Another continuation of this work is to adopt CLIR for in-context learning with LLMs, which are currently instruction-tuned with parallel examples, based on source-side similarities Moslem et al.
-(2023).
-Fine-tuning LLMs with just target side examples retrieved with CLIR is one priority of our future work.
-Admittedly, we worked with high ressource languages to assess the viability of our method.
-Considering a low ressource setup, provided large enough monolingual corpora are available, is a natural use case when it is even harder to access to parallel data.
-Such experiments can further assess the strenght of our method.
+Limitations
+In this work, we have focused on extending a translation setting that is common in the translation industry, where translation memories have repeatedly been shown to speed up the translation process and improve the consistency of the resulting texts.
+In such setting, it is also common to have access to large monolingual resources, as has been studied eg., by (Tezcan et al., 2024) for the legal domain and for (Tamura et al., 2023) in the scientific domain.
+By design, this setting is only applicable for so-called “high-resource" language pairs, in which high-quality systems can be built and improved upon using auxiliary resources.
+In our selection of language pairs, we have favored domain diversity over language diversity, to better highlight how the availability of relevant examples impacts the overall translation quality.
+As our experiments show, the effectiveness of the cross-lingual approach does not depend on the similarity of the source and target languages, but mostly relies on the quality of cross-lingual sentence alignment.
+As shown e.g., by (Feng et al., 2022), very precise alignments can be achieved, even for distant language pairs such as English-Chinese.
+We have thus deliberately chosen to exclude translation from and into “low-resource” languages, for which many resources and components are still difficult to collect (eg., parallel corpora to train sufficiently good NMT baselines, to be used for back translation; and also training and test data in specialized domains, where the TM-based approach is most likely to retrieve very good matches).
+We reckon that also improving generic NMT systems that could handle translations for those language pairs is another important research area, albeit distinct from what we have chosen to study in this work.
+Ethical Statement
+There are no ethical issues.
+Acknowledgments
+This research was funded by the French Agence Nationale de la Recherche (ANR) under the project TraLaLaM (“ANR-23-IAS1-0006”).
+This work was performed using HPC resources from GENCI–IDRIS (Grant 2024-A0161015117).
 Appendix A Parallel Data Specification
@@
 The en-fr data is a clean version [^12] We filter noisy parallel pairs with COMETKiwi Rei et al.
-[2022] and prepare a new train/valid/test split.
-of the one used by [Xu et al., 2020].
+(2022) and prepare a new train/valid/test split.
+of the one used by (Xu et al., 2020).
 Validation and test sets each contain $1,000$ segments.
-The de-en corpus is borrowed unchanged from [Aharoni and Goldberg, 2020] and is also used in [Cai et al., 2021; Agrawal et al., 2023; Bouthors et al., 2024].
+The de-en corpus is borrowed unchanged from (Aharoni and Goldberg, 2020) and is also used in (Cai et al., 2021; Agrawal et al., 2023; Bouthors et al., 2024).
 The associated validation and test sets contain $2,000$ segments.
+The per-domain corpus sizes are available in table 8.
+The en-uk corpus is the one provided by (Tezcan et al., 2024), containing about 286K training sentences, 2000 validation sentences and 1898 test sentences.
+It corresponds to various legal texts.
+Note that the preprocessing stages were performed with subword-nmt (Sennrich et al., 2016b) for tokenization and BPE, and fairseq (Ott et al., 2019) for binarization.
 Appendix B Search Parameterization
-For both methods, we search up to $k$=3 examples.
+For both methods, we search up to kk=3 examples.
 B.1 Fuzzy Matching
-The execution of the fuzzy matching methods (fuzzy-src, fuzzy-bt and fuzzy-gold) is done via an open source library [^9] https://github.com/SYSTRAN/fuzzy-match.
+The execution of the fuzzy matching methods (fuzzy-src, fuzzy-bt and fuzzy-gold) is performed using an open source library [^13] https://github.com/SYSTRAN/fuzzy-match.
 During the evaluation, there is no filter applied before scoring the segments with $\operatorname{Lev}$.
 However, the examples associated to the training samples are selected with $\operatorname{Lev}$and a BM25 filter.
-This ensures a reasonable time to build the set of examples for traning.
+This ensures a reasonable time to build the set of examples for training.
 B.2 FAISS
-The execution of the CLIR techniques is done with library FAISS [Douze et al., 2024].
-The $k$NN search is performed on GPU (IndexFlatIP), without IVF (for faster search) or quantizer (for lower memory footprint) to ensure the retrieval of the best matches.
+The execution of the CLIR techniques is done with library FAISS (Douze et al., 2024).
+The kkNN search is performed on GPU (IndexFlatIP), without IVF (for faster search) or quantizer (for lower memory footprint) to ensure the retrieval of the best matches.
 Indeed, both optimizations can prevent the model from finding the optimal $k$ nearest neighbors.
 Appendix C NMT Architectures Configuration
-Both TM3-LevT and NFA models rely on the default backbone Transformer architecture [Vaswani et al., 2017].
+Both TM3-LevT and NFA models rely on the default backbone Transformer architecture (Vaswani et al., 2017).
 We train one instance of TM3-LevT for each language pair (en-fr and de-en), using only the training data available in the corresponding datasets.
@@
 As we compare retrieval techniques for fixed MT architectures, these overlaps do not affect on our main conclusions (see also the discussion in §6.2).
-In addition, we use the multilingual NLLB model [Costa-jussà et al., 2024] out-of-the-box for both language pairs.101010version: NLLB-200-distilled-1.3B
+In addition, we use the multilingual NLLB model (Costa-jussà et al., 2024) out-of-the-box for both language pairs.141414version: NLLB-200-distilled-1.3B
 https://huggingface.co/facebook/nllb-200-1.3B
@@
 Contamination is discussed in §6.2.
-Appendix D Computational Cost of Retrieval
+Appendix D Adapting EuroLLM
+LLMs are well suited for in-context learning in $k$-shot settings (Radford et al., 2019): a set of $k$ demonstrations of the task are presented to the model so that the completion of the task fulfills the prompt pattern/template.
+In our case, the situation is slightly different as we only include target-side examples, which are not suitable demonstrations of the translation task.
+To adapt an LLM to this setting, we first compare (a) prompting and (b) fine-tuning strategies on the Wikipedia domain.
+[^15] Another alternative, that we do not consider in this work, would be to use backtranslation to automatically craft the missing source for in-context examples.
+Regarding (a), we try to take advantage of the in-context abilities of the Instruct version of EuroLLM-9B, simulating in-context examples with the follow prompt patterns:
+  • naive:
+    Consider the following French similar translations.
+    {k examples in French}
+    Translate the following text from English to French.
+    English:
+    {source text}
+    French:
+  • empty:
+    Consider the following English-French examples.
+    English:
+    French:
+    {French example #1}
+    Translate the following text from English to French.
+    English:
+    {source}
+    French:
+  • copy:
+    Consider the following English-French examples.
+    English:
+    {French example #1}
+    French:
+    {French example #1}
+    Translate the following text from English to French.
+    English:
+    {source}
+    French:
+The naive prompt gives a simple description of the task, where only the target side of examples are introduced.
+The copy and empty prompts simulate a standard $k$-shot prompt, where the missing source text is either left empty or contains a copy of the target text.
+We compare these prompts with a 0-shot setting, where no example is provided to the model (starting directly with "Translate the following text" for any of the three templates).
+Regarding (b), we fine-tune the base EuroLLM-9B model with the naive prompt with up to 20K sentences from the training set for each domain.
+Fine-tuning is performed with QLoRA (Dettmers et al., 2023) with a rr=16, $\alpha$=32, dropout=0.05, no bias, applied on all linear layers, and 4-bit quantization.
+Training is performed with huggingface implementation [^16] https://huggingface.co/docs/transformers/trainer with a batch size of 32 for a single epoch and a learning rate of 2e-10.
+The in-context examples used for training are retrieved using the fuzzy-src strategy.
+Instruct ft-base
+0-shot naive copy empty naive
+(k=0) (k=3) (k=3) (k=3) (k=3)
+36.5 36.5 38.4 37.5 49.5
+Table 9: BLEU score of various EuroLLM-9B based systems on the Wikipedia test set (0-shot and 33-shot). Target examples are retrieved with LaBSE with a custom threshold of $0.6$.
+We observe the superiority of the fine-tuning strategy compared to the various prompt-based approaches in Table 9.
+As some of these differences might be explained by the implicit language and domain adaptation that takes place during model fine-tuning, we additionally fine-tune the base model with a standard $k$-shot prompt, including the source and target side of each examples.
+This setting directly compares with the instruct model, the only difference being the adaptation that is performed during fine-tuning; it also allows us to assess the performance loss that happens when using only target side examples.
+Results are in Table 10 for a mixture of $5$ domains.
+EuroLLM model source k BLEU
+Instruct - 0 42.9
+3 54.0
+ft-base (src+tgt) - 0 46.7
+3 58.5
+ft-base (tgt) - 0 47.3
+3 57.5
+Table 10: Average BLEU score of various EuroLLM-9B system fine-tuning schemes for ECB, EMEA, Europarl, JRC-Acquis and Wikipedia test sets (k-shot).
+Parallel examples are retrieved with fuzzy-src using a custom threshold of 0.
+In Table 10, we observe that in fact the domain/language adaptation of EuroLLM accounts for about +4 BLEU points (ft-base (src+tgt) vs.
+instruct).
+For the sole Wikipedia domain, this adapation yields about +6 BLEU points improvement.
+This confirms that method (b) is in fact clearly superior to simple prompting strategies.
+Therefore, this is the method used for all the results reported in Section 6.
+Moreover, we observe that the absence of the source side of $k$-shot examples is slightly detrimental to the model performance ($\approx$ -1 BLEU point between the two fine-tuned versions, both for the 0-shot and $3$-shot settings), suggesting that the target side of the retrieved examples contains most of the information needed to guide the model towards the desired translation.
+Appendix E Computational Cost of Retrieval
 The computational cost can be dived into two categories: a fixed cost (independant from the number of sentences to translate) and a variable cost (proportional to this number).
 The retrieval process for training falls into the former, as well as the encoding of the whole set of monolingual target segments (for CLIR techniques) and the indexation of the TM (for fuzzy matching techniques).
-The variable costs lie in the fuzzy match search with filtering and the computation of $\operatorname{Lev}$, and in the CLIR search with the encoding of the source sentences, then the $k$NN search.
+The variable costs lie in the fuzzy match search with filtering and the computation of $\operatorname{Lev}$, and in the CLIR search with the encoding of the source sentences, then the kkNN search.
 In our experiments, we removed the filter (BM25) so that we obtain the best possible fuzzy matches (fuzzy-src, fuzzy-bt and fuzzy-gold), at the cost of a $\sim$100 times higher latency.
@@
   • the fuzzy matching is performed on a 8 cores CPU, thus parallelizing the search on 8 simultaneous source sentences;
-  • as for CLIR, source sentences are encoded in batches of size 50, and the $k$NN search is handled by library FAISS. We used a V100-32GB GPU.
+  • as for CLIR, source sentences are encoded in batches of size 50, and the kkNN search is handled by the FAISS library. We used a V100-32GB GPU.
+Appendix F Retrieval and Translation Examples
+retriever $\operatorname{Lev}$
+reference Le mannequin apparaît dans plusieurs séries télévisées au début des années 2000. (The model also appears in several TV shows in the early 2000s.)
+train LaBSE 0.14 Cette série fut diffusée sur TF1 pendant l’été 2000. (This show was broadcast on TF1 during the summer of 2000.)
+train ft-LaBSE-MAE 0.23 Elle est aussi apparue dans des séries télévisées. (She also played in some TV shows.)
+mono LaBSE 0.46 Elle se fait connaître par sa série au début des années 2000. (She made a name for herself thanks to her show in the early 2000s.)
+mono ft-LaBSE-MAE 0.62 Elle débute dans des séries télévisées au début des années 1990. (She made her debut in some TV Shows in the early 1990s.)
+reference Shasta a connu une histoire explosive et éruptive. (Shasta has had an explosive and eruptive history.)
+train LaBSE 0.06 L’histoire de la région du Huaynaputina est marquée par un magma riche en silice. (The history of the Huaynaputina region is marked by silica-rich magma. )
+train ft-LaBSE-MAE 0.44 Elle a une histoire très singulière. (She has a very peculiar history.)
+mono LaBSE 0.07 Il s’agit d’un des cônes volcaniques du mont Shasta. (This is one one the vulcanic cones of Mount Shasta.)
+mono ft-LaBSE-MAE 0.56 Elle a connu une histoire riche en rebondissements. (Its history is full of twists and turns.)
+Table 11: Two illustrations of retrieved examples for the Wikipedia test set w.r.t.
+the chosen cross-lingual retriever, and the retrieval pool (train set vs.
+monolingual corpus).
+Exact matches are marked in bold.
+Indicative translations of each French segment into English are also provided.
+An illustration of the retrieved examples is provided is table 11.
+It highlights the enhanced capability of ft-LaBSE-MAE to retrieve lexically closer examples, rather than semantically close ones.
+This is particularly clear in the second example, which mentions Shasta, a volcano.
+While LaBSE retrieves topic-related segments, ft-LaBSE-MAE tends to select off-topic sentences that nonetheless contain a higher number of shared lexical items.
+Appendix G Back-Translation
+Back-translation (BT) has long been identified as a very effective technique to handle monolingual data in Statistical Machine-Translation (Bojar and Tamchyna, 2011), then in Neural Machine Translation (Sennrich et al., 2016a; Currey et al., 2017; Edunov et al., 2018; Burlot and Yvon, 2018; Caswell et al., 2019).
+Assuming the task is to translate from L1 into L2, and that monolingual texts in L2 are available, the BT approach translates these “backward” into L1, to obtain an artificial parallel corpus, pairing automatically generated source sentences with actual human-written target sentences.
+In NMT, it is custom to jointly use “natural” and “artificial” subsets of data to train and/or adapt translation models.
+This is illustrated in the first two columns (a) and (b) in Table 12, using no-retrieval ($k=0$), and corresponding respectively to the default (only parallel data) and BT-augmented settings.
+[^17] We omit the case where only artificial data is used, as in fully unsupervised machine translation (Lample et al., 2018).
+MT setting (a) (b) (c) (d) (e) (f)
+Parallel data
+train ($k=0$) - - - -
+train ($k\geq 0$) - -
+test ($k\geq 0$) - -
+BT-ed data
+train ($k=0$) - - - - -
+train ($k\geq 0$) - - - -
+test ($k\geq 0$) - - - -
+Table 12: Using back-translation in NMT and RANMT
+Column (c) in Table 12 illustrates standard RANMT, where $k\geq 0$ parallel examples are used during training and at inference; columns (d), (e) and (f) correspond to three possible ways to integrate artificial data: (d) only during training, and retrieve selectively from the high-quality parallel data; (e) only during inference, extending a generic RANMT with relevant data; (f) both during training and at inference as in Tezcan et al.
+(2024).
+Our use of back-translation in the experiments of Section 5 corresponds to column (e), a setting we deem representative of typical use of monolingual data in actual applications, and which crucially does not require retraining stage.
+Note that the same scenario is used in our CLIR approach, which, in addition, fully dispenses with the back-translation stage.
+Appendix H Experiments in English-Ukrainian
+To further emphasize the ability of our method to retrieve relevant examples even in the absence of direct lexical overlap between source and target segments, we experiment with the translation from English into Ukrainian, with the source using the Latin script, and the target the Cyrillic script.
+We used the same dataset as Tezcan et al.
+(2024), corresponding to legal texts.
+The parallel corpus is split into train, validation and test sets, each containing respectively 286,417, 2000 and 1898 sentences.
+This corpus is augmented with a monolingual in-domain dataset of 1,461,320 Ukrainian sentences.
+Our experiments are conducted with the same NFA architecture as the experiments in Section 5, as it produces translations with higher estimated-quality than the non-autoregressive system.
+This NFA model has been trained on 5M en-uk parallel pairs from various domains, including the in-domain train set.
+Retrieval method Corpus BLEU COMET
+no example - 51.2 92.7
+fuzzy-src train 54.2 92.8
+LaBSE train 53.9 92.7
+ft-LaBSE-MSE train 54.3 92.8
+LaBSE train+mono 57.2 92.9
+ft-LaBSE-MSE train+mono 57.7 93.0
+Table 13: BLEU and COMET scores obtained by the NFA model according to the retrieval method. Significance w.r.t. to the LaBSE counterpart is indicated by .
+We fine-tune LaBSE using the proposed MSE lexical loss, with examples retrieved via fuzzy-src from the training set.
+We then evaluate and compare the performance of fuzzy-src, LaBSE, and ft-LaBSE-MSE by computing BLEU and COMET scores of translations produced by the NFA model.
+For LaBSE and ft-LaBSE-MSE, we consider two retrieval settings — train and train+mono — which correspond to the datasets from which target-side examples are retrieved for the CLIR approach.
+Results are in Table 13, scores are computed respectively by SacreBLEU and COMET-22, significance tests also use the SacreBLEU implementations with paired bootstrap resampling (n=1000, p=0.05).
+We can make the following observations: (a) the baseline NFA model is about 3 BLEU points better than baseline encoder-decoder model; (b) CLIR here again matches the default RAMT set-up, where retrieval only exploits the parallel training data; (c) augmenting the retrieval pool with monolingual data again induces a significant performance boost (+3.4 BLEU), when retrieval relies on the fine-tuned version of LaBSE.
+These observations are consistent with what is reported in the main text.
+Appendix I COMET Scores
+For the full picture, we also report the COMET scores of each domain for all our experiments in Tables 14, 15 and 16.
+English-French German-English
+ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
+fuzzy-gold 70.5 85.0 45.8 73.2 71.9 58.9 24.3 35.9 21.2 52.1 29.8 30.3 -98.3 19.9 29.7 -24.8
+fuzzy-src 69.0 81.5 44.7 67.6 69.8 47.9 24.7 33.5 19.5 49.2 25.0 24.1 -85.8 16.6 21.8 -28.2
+fuzzy-bt 61.6 75.0 45.6 55.5 65.8 37.3 24.1 30.1 17.7 45.4 24.2 11.4 -110.6 -11.0 21.3 -25.4
+LASER 65.8 76.3 45.1 58.1 66.0 34.7 23.2 29.2 19.9 45.0 24.5 18.5 -73.0 19.7 27.5 -31.0
+LaBSE 65.5 77.2 42.3 62.9 67.9 42.8 23.3 29.7 19.0 47.1 23.3 21.2 -86.4 21.4 26.3 -30.1
+ft-LaBSE-MSE 68.0 81.1 44.4 66.5 69.8 48.0 24.4 31.9 18.5 49.6 23.8 18.5 -76.3 8.7 15.6 -29.1
+ft-LaBSE-MAE 69.0 80.7 44.2 65.5 68.4 47.0 23.8 32.7 18.9 49.8 24.9 20.2 -70.6 14.1 19.1 -28.6
+ft-LaBSE-Rank 67.3 79.6 44.8 63.4 68.8 46.5 23.8 32.2 19.7 47.9 24.5 20.7 -84.9 14.0 17.6 -28.5
+Table 14: Per domain COMET scores for TM3-LevT.
+English-French German-English
+ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
+no example 58.9 55.9 39.9 48.8 62.3 42.8 50.1 45.6 43.9 43.5 36.4 82.2 68.9 83.2 84.7 80.2
+fuzzy-gold 90.5 90.9 87.9 89.8 90.7 85.8 88.2 84.6 85.4 86.6 85.2 84.5 71.4 88.2 86.3 80.6
+fuzzy-src 90.4 90.5 87.8 89.3 90.5 85.2 88.2 84.5 85.4 86.3 85.0 84.1 70.0 87.9 85.9 80.4
+fuzzy-bt 90.0 90.1 87.7 88.6 89.9 84.4 88.2 84.3 85.2 86.0 84.9 83.5 70.1 87.5 85.2 80.3
+LASER 90.0 90.1 87.8 88.8 90.0 84.4 88.2 84.2 85.4 86.4 84.7 83.5 70.0 87.4 85.3 80.4
+LaBSE 90.0 90.3 87.5 89.0 89.8 84.7 88.2 84.1 85.3 86.4 84.9 83.8 70.3 87.2 85.5 80.3
+ft-LaBSE-MSE 90.3 90.5 87.7 89.2 90.4 85.1 88.2 84.5 85.3 86.5 84.8 83.8 70.3 87.8 85.6 80.4
+ft-LaBSE-MAE 90.3 90.5 87.7 89.2 90.4 85.1 88.2 84.5 85.4 86.4 84.8 83.7 70.2 87.7 85.7 80.3
+ft-LaBSE-Rank 90.2 90.4 87.7 89.2 90.5 85.3 88.2 84.5 85.3 86.5 84.9 83.8 70.5 87.8 85.7 80.4
+Table 15: Per domain COMET scores for NFA.
+English-French German-English
+ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
+no example 88.2 88.5 87.7 85.3 89.1 82.8 86.2 84.3 85.0 85.8 86.3 85.9 74.7 86.9 85.1 80.5
+fuzzy-gold 90.2 90.9 87.7 90.1 90.3 87.5 86.3 85.2 85.4 89.0 87.0 88.0 75.5 88.4 87.0 80.8
+fuzzy-src 90.0 90.5 87.8 89.3 90.1 86.2 86.2 85.1 85.3 88.4 86.8 86.5 74.9 87.2 85.5 80.7
+fuzzy-bt 88.6 89.0 87.7 87.2 89.2 84.0 86.2 84.7 85.3 87.6 86.6 86.0 74.9 87.2 85.6 80.6
+LASER 89.8 90.0 87.7 88.1 90.0 84.6 86.3 84.9 85.2 88.3 86.6 86.3 75.0 87.9 85.9 80.6
+LaBSE 90.0 90.2 87.7 88.6 90.1 85.8 86.3 84.9 85.2 88.2 86.6 86.5 75.0 87.9 86.1 80.5
+ft-LaBSE-MSE 90.0 90.5 87.7 88.9 90.1 86.4 86.3 85.1 85.3 88.5 86.8 86.5 75.1 87.8 85.8 80.6
+ft-LaBSE-MAE 89.9 90.6 87.7 89.1 90.1 86.4 86.2 85.0 85.2 88.4 86.7 86.7 75.1 87.9 86.0 80.5
+ft-LaBSE-Rank 89.8 90.3 87.7 89.2 90.0 86.2 86.3 84.9 85.3 88.6 86.6 86.6 75.0 87.8 86.0 80.5
+Table 16: Per domain COMET scores for EuroLLM.